### 20 Mar 2026 v7

# Learnability Window in Gated Recurrent Neural Networks

Abstract
We develop a statistical theory of temporal learnability in recurrent neural networks, quantifying the maximal temporal horizon $\mathcal{H}_{N}$ over which gradient-based learning can recover lag-dependent structure at finite sample size $N$. The theory is built on the effective learning rate envelope $f(\ell)$, a functional that captures how gating mechanisms and adaptive optimizers jointly shape the coupling between state-space transport and parameter updates during Backpropagation Through Time. Under heavy-tailed ($\alpha$-stable) gradient noise, where empirical averages concentrate at rate $N^{-1/\kappa_{\alpha}}$ with $\kappa_{\alpha}=\alpha/(\alpha-1)$, the interplay between envelope decay and statistical concentration yields explicit scaling laws for the growth of $\mathcal{H}_{N}$: logarithmic, polynomial, and exponential temporal learning regimes emerge according to the decay law of $f(\ell)$. These results identify the envelope decay geometry as the key determinant of temporal learnability: slower attenuation of $f(\ell)$ enlarges the learnability window $\mathcal{H}_{N}$, while heavy-tailed gradient noise compresses temporal horizons by weakening statistical concentration. Experiments across multiple gated architectures and optimizers corroborate these structural predictions.
1 Introduction
Recurrent neural networks (RNNs) are fundamental models for processing sequential data, yet their ability to learn long-range temporal dependencies remains only partially understood. Gated architectures such as the LSTM and GRU have dramatically improved numerical stability and empirical performance, but it is still unclear which temporal dependencies are statistically recoverable under finite data. Most existing analyses focus on dynamical stability, spectral properties of Jacobian products, or mean-field approximations. These approaches characterize when gradients do not explode or vanish, but they do not provide a statistical criterion for when transported gradient signals remain distinguishable from noise during training. In practice, even stable gradients may be too attenuated or too noisy to carry usable information. Thus, stability alone does not determine temporal learnability.
Our previous work showed that gating induces heterogeneous time scales that shape both state evolution and gradient transport [38]. Here we move beyond purely dynamical analyses and develop a statistical theory of finite-horizon learnability. The central objects are the effective learning rates $\mu_{t,\ell}$, which quantify how Backpropagation Through Time (BPTT) reweights gradient signals across temporal lags for each neuron. These quantities arise from the interaction between gate-induced transport of gradients through recurrent dynamics and the base learning rates imposed by the optimizer. Their aggregate magnitude defines the effective learning rate envelope $f(\ell)=\|\mu_{t,\ell}\|_{1}$, which governs the effective strength of lagged gradient contributions to parameter updates during training. Although effective learning rates were originally [38] derived under SGD with a fixed global learning rate $\mu$, we generalize them here to adaptive optimizers such as Adam, where each parameter receives its own time-varying rate. The generalization replaces the global $\mu$ with a neuron-specific adaptive base rate $\Lambda^{(q)}_{r,\ell}$, obtained by projecting the parameter-space preconditioner onto each neuron via a Rayleigh-quotient construction, while leaving the gate-induced transport factors unchanged.
The central question addressed in this paper is: given a finite number of training sequences, up to which temporal horizon can dependencies be statistically detected? To answer this, we introduce the learnability window $\mathcal{H}_{N}$, defined as the largest lag for which gradient information remains recoverable at sample size $N$. Under heavy-tailed ($\alpha$-stable) gradient noise, empirical averages concentrate at rate $N^{-1/\kappa_{\alpha}}$ with $\kappa_{\alpha}=\alpha/(\alpha-1)$. We show that the interaction between this concentration rate and the decay geometry of the envelope $f(\ell)$ determines the scaling behavior of $\mathcal{H}_{N}$. In particular, exponential, polynomial, and logarithmic attenuation of $f(\ell)$ induce qualitatively distinct growth laws for the learnability window, providing a universal classification of temporal learning regimes.
Although the detectability relation analytically separates two factors, the decay geometry of the envelope and the statistical concentration of gradient noise, these mechanisms are not independent during training. The temporal transport of gradients is shaped jointly by gating dynamics and optimizer adaptation, while the statistics of gradient fluctuations evolve with the optimization process. Empirically, regimes exhibiting slow envelope decay often co-occur with persistent heavy-tailed gradient fluctuations, reflecting the coupled evolution of state and parameter dynamics.
Importantly, heavy-tailed noise does not enhance learnability. Instead, it weakens statistical concentration and therefore raises the data requirements for reliable detection of long-range dependencies. This interplay is reflected in the relation $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, which links envelope decay to the sample complexity of temporal credit assignment. Temporal learnability is thus determined by the balance between envelope attenuation and noise concentration: heterogeneous time-scale structure tends to slow the decay of $f(\ell)$ and extend the learnability window, whereas heavy-tailed gradient noise counteracts this effect by reducing statistical efficiency.
Contributions.
Relative to prior dynamical analyses of gated RNNs, this paper makes four contributions. (i) We formalize the learnability window $\mathcal{H}_{N}$ as a finite-sample measure of recoverable temporal dependencies and incorporate heavy-tailed ($\alpha$-stable) gradient noise into the analysis of recurrent training. (ii) We derive explicit scaling laws that classify temporal learning regimes into canonical classes, linking envelope decay $f(\ell)$, sample-complexity requirements $N(\ell)$, and the growth of the learnability window $\mathcal{H}_{N}$. (iii) We generalize the effective learning rate framework from SGD to adaptive optimizers by introducing a neuron–lag projection of the parameter-space preconditioner. (iv) We empirically validate these structural predictions across multiple gated architectures.
Paper organization.
The remainder of the paper is organized as follows. Section 2 reviews related literature. Section 3 recalls the BPTT framework under both SGD and adaptive optimizers. Section 4 introduces the generalized effective learning rate formulation used throughout the paper. Architecture-specific Jacobian derivations and transport factors for all recurrent models are provided in Appendix B. Section 5 develops the theoretical framework of the learnability window. Section 6 presents empirical validation. Section 7 concludes with implications and future directions. Additional technical derivations and supporting analyses are provided in the appendices (Appendices AK).
2 Related work
The introduction of gating mechanisms in LSTM [20] and GRU [9] was pivotal for controlling temporal credit assignment in recurrent neural networks. Early theoretical works linked the forget gate to an interpretable exponential decay and controllable memory retention [15, 51], while more recent studies connect gating to continuous-time or implicit ODE perspectives [5, 16]. These contributions illuminate how gates shape state dynamics, but do not analyze how gating multiplicative structures affect parameter updates and the capacity for learning long-range dependencies. Our framework bridges that gap by deriving per-lag effective learning rates from products of Jacobians and linking them to finite-horizon learnability.
A complementary line of work models gates as learned rescaling or time-warping mechanisms. Tallec and Ollivier [51] formalized gating as adaptive time dilation, and recent geometry-based analyses [33] study how gating choices influence information flow. We build on these ideas by explicitly casting gate-derived Jacobian terms into effective learning rates and embedding them in detectability bounds via mutual information / Fano theory.
The classical vanishing/exploding gradients problem motivated many stabilization techniques: clipping, spectral regularization, and orthogonality or unitary RNNs [3, 43, 1, 54, 26]. Hierarchical, dilated, or skip architectures shorten propagation paths [30, 6, 10], while continuous-time and neural ODE approaches enhance stability and expressivity [45, 18]. Works on orthogonal / dynamical isometry control the conditioning of gradient transport [7, 44] but do not address whether transported gradients still carry statistically reliable information over long lags. In contrast, we analyze statistical detectability, showing that even when products of Jacobians remain numerically stable, learning can be ineffective if the envelope decays too rapidly. Related difficulties have long been observed in tasks with delayed supervision or rewards [3], where the learning signal must be propagated across long temporal gaps; however, existing analyses typically focus on optimization stability rather than statistical detectability across lags.
A related but distinct line of work concerns memory capacity in reservoir computing and echo state networks, where the recurrent dynamics are kept fixed and only a readout layer is trained [53, 24, 13]. In this setting, classical results characterize the ability of a randomly driven recurrent system to retain information about past inputs over time, typically quantified via linear or nonlinear memory capacity measures. Such analyses focus on expressivity and information retention properties of untrained dynamical systems, rather than on the optimization dynamics of recurrent weights under gradient-based learning. By contrast, the present work addresses a complementary question: how training dynamics themselves impose horizon-dependent limits on learnability through the interaction between temporal Jacobian transport, effective learning rates, and noise. In particular, our framework does not rely on fixed reservoirs, but instead characterizes how credit assignment and detectability degrade under heavy-tailed gradient noise, limiting the realizable horizon even in architectures with sufficient representational capacity.
In the broader learning systems literature, trainability has been studied via mean-field theory, spectral initialization, curvature, and the geometry of parameter landscapes [47, 44, 35, 11, 40, 36, 48, 56, 57, 4, 28]. In recurrent and transformer settings, recent work further links trainability to Jacobian spectra, Fisher information, and anisotropic learning rates [55, 2, 50, 14]. Related analyses study gradient signal-to-noise ratios and noise scales as a function of batch size, curvature, and optimizer dynamics, with the aim of understanding optimization efficiency and generalization behavior [41, 25]. Zhou et al. [58] investigate the generalization gap between stochastic gradient descent (SGD) and Adam, attributing Adam’s performance to a sharpness reduction effect caused by adaptive coordinate-wise scaling and noting an inherent clipping behavior. However, these approaches typically analyze state dynamics and parameter dynamics in isolation. They do not treat learning as a coupled dynamical system in which temporal transport through recurrent Jacobians interacts with stochastic optimization noise, nor do they characterize the resulting universal scaling regimes governing temporal learnability and the sample complexity of detecting long-range dependencies.
Simsekli et al. [49] show that mini-batch SGD noise often exhibits $\alpha$-stable, heavy-tailed behavior rather than Gaussian tails. Subsequent research has further examined the algorithmic and theoretical consequences of such heavy-tailed gradient statistics. Hübler et al. [22] establish a unified framework connecting gradient clipping and normalization, proving that both act as implicit variance control mechanisms for $\alpha$-stable noise and can stabilize training without sacrificing convergence speed. Finally, Liu and He [37] extend these ideas to online convex optimization with heavy-tailed stochastic gradients, providing convergence guarantees and adaptive algorithms that remain stable even when gradient moments of order two do not exist. In our work, we adopt the empirical reality of heavy-tailed gradient noise as a basic modeling premise.
Finally, structured state-space models achieve long-range dependencies via implicit recurrence or convolutional kernels [16, 17]. Our focus remains on classical recurrent systems trained by BPTT: we deliver measurable per-lag detection thresholds and learnability windows backed by heavy-tailed gradient statistics.
3 Backpropagation through time
Training RNNs follows the same fundamental principle as feedforward models: parameter updates are obtained through SGD or variations thereof [46]. Given trainable parameters $\theta$ and learning rate $\mu$, the SGD update at iteration $r$ is
$\theta_{r+1}\;=\;\theta_{r}-\mu\,\nabla_{\theta}\mathcal{L}(\theta_{r}),\qquad\mathcal{L}\;=\;\sum_{t=1}^{T}\mathcal{E}_{t},$ (1)
where $\mathcal{E}_{t}$ denotes the instantaneous loss at time $t$ within a sequence of length $T$.
Under an adaptive optimizer (e.g., Adam, AdamW, RMSprop), the update can be written as a diagonally preconditioned gradient descent step [28, 46]
$\theta_{r+1}=\theta_{r}-\Lambda_{r}\nabla_{\theta}\mathcal{L}(\theta_{r}),\qquad\Lambda_{r}=\mathrm{diag}(\lambda_{1,r},\dots,\lambda_{P,r})\in\mathbb{R}^{P\times P},$ (2)
where $P=\dim\theta$ and $\lambda_{i,r}>0$ are per-parameter adaptive learning rates determined by the optimizer state. For example, in Adam-type methods, these rates depend on running estimates of gradient moments.
In recurrent architectures, computing the gradient $\nabla_{\theta}\mathcal{L}$ requires unrolling the network dynamics through time and accounting for how earlier states influence later losses–a process known as Backpropagation Through Time.
Let $h_{t}$ denote the recurrent state at time $t$. The total gradient of the loss with respect to the parameters can be written as
$\nabla_{\theta}\mathcal{L}\;=\;\sum_{t=1}^{T}\frac{\partial\mathcal{E}_{t}}{\partial\theta}\;=\;\sum_{t=1}^{T}\frac{\partial\mathcal{E}_{t}}{\partial h_{t}}\sum_{\ell=1}^{t}\frac{\partial h_{t}}{\partial h_{\ell}}\frac{\partial h_{\ell}}{\partial\theta}.$ (3)
The outer sum aggregates contributions from all time steps, while the inner chain rule expresses how parameter perturbations at earlier times influence the current loss through the recurrent dynamics. To make this dependence explicit, define $J_{j}\;=\;\frac{\partial h_{j}}{\partial h_{j-1}}$ and $B_{\ell}(\theta)\;=\;\frac{\partial h_{\ell}}{\partial\theta}$. Here $J_{j}$ is the state Jacobian, which quantifies how the state evolves in response to infinitesimal perturbations of the previous state, and $B_{\ell}(\theta)$ is the parameter-state Jacobian, measuring the instantaneous sensitivity of the state to the parameters at step $\ell$. Substituting these definitions into Eq. (3), the gradient contribution of a specific loss term $\mathcal{E}_{t}$ becomes
$\frac{\partial\mathcal{E}_{t}}{\partial\theta}\;=\;\delta_{t}^{\top}\sum_{\ell=1}^{t}\mathcal{M}_{t,\ell}\,B_{\ell}(\theta),\qquad\mathcal{M}_{t,\ell}\;=\;\prod_{j=\ell+1}^{t}J_{j},$ (4)
where $\delta_{t}=\partial\mathcal{E}_{t}/\partial h_{t}$ denotes the local loss gradient at time $t$. The matrix product $\mathcal{M}_{t,\ell}$, often referred to as the state-transition Jacobian product, transports this signal backward through the sequence, modulating both its magnitude and direction as it interacts with the intermediate state Jacobians $J_{j}$.
4 Generalized effective learning rates under adaptive optimization
This section extends the effective learning rate framework to adaptive optimizers. Appendix B derives the neuronwise effective learning rates obtained from the first-order expansion of recurrent Jacobian products for all architectures considered in this paper. These derivations show that the lag-dependent learning rates induced by BPTT can be expressed as the product of two components: a scalar optimizer step size and a transport factor determined by the recurrent dynamics.
To streamline notation, we collect the diagonal contributions of the Jacobian-product expansion into the transport factor $\Gamma^{(q)}_{t,\ell}$, which aggregates the zeroth- and first-order terms arising from the expansion. For example, in the LSTM this factor takes the form
$\Gamma^{(q)}_{t,\ell}=\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}$,
corresponding to the sum of the zeroth- and first-order transport contributions along the cell pathway.
Under plain SGD, the optimizer applies a uniform base learning rate $\mu$ to all parameters. The neuronwise effective learning rates therefore factorize as
$\mu^{(q)}_{t,\ell}=\mu\,\Gamma^{(q)}_{t,\ell}$, (5)
where $\Gamma^{(q)}_{t,\ell}$ depends only on the recurrent Jacobian transport and thus entirely determines the decay geometry of the effective learning rates across temporal lags.
Modern recurrent networks, however, are typically trained using adaptive optimizers such as Adam or RMSprop, which apply a diagonal preconditioner to the gradient. In this setting, the global base rate $\mu$ is replaced by a neuron-specific adaptive base rate $\Lambda^{(q)}_{r,\ell}$ derived from the optimizer preconditioner. The resulting generalized effective learning rates take the form
$\mu^{(q)}_{t,\ell}=\Lambda^{(q)}_{r,\ell}\Gamma^{(q)}_{t,\ell}$. (6)
The adaptive base rate $\Lambda^{(q)}_{r,\ell}$ is obtained by projecting the diagonal optimizer preconditioner onto the parameter-space direction associated with neuron $q$ at lag $\ell$. This projection is given by the Rayleigh quotient
$\Lambda^{(q)}_{r,\ell}=\frac{B_{\ell}^{(q)}\,\Lambda_{r}\,(B_{\ell}^{(q)})^{\top}}{B_{\ell}^{(q)}(B_{\ell}^{(q)})^{\top}},$ (7)
where $B_{\ell}^{(q)}$ denotes the parameter-state sensitivity of neuron $q$ and $\Lambda_{r}$ is the diagonal optimizer preconditioner. A geometric interpretation and derivation of this projection are provided in Appendix D.
5 Learnability under heavy-tailed gradient noise
[…]
5.1 The envelope
The precise form of the effective learning rates depends on the optimizer (Section 4). However, the learnability analysis depends on the effective learning rates $\mu^{(q)}_{t,\ell}$ only through their aggregate magnitude across neurons. We therefore introduce the envelope
[…]
This scalar summarizes the total strength with which lag-$\ell$ gradient contributions are amplified or attenuated by the combined action of recurrent transport and optimizer scaling.
The envelope admits the following decomposition
$f(\ell)=\underbrace{\mu\sum_{q=1}^{H}|\Gamma^{(q)}_{t,\ell}|}_{f_{\mathrm{gates}}(\ell)}+\underbrace{\sum_{q=1}^{H}\Delta\Lambda^{(q)}_{r,\ell}\,|\Gamma^{(q)}_{t,\ell}|}_{f_{\mathrm{adapt}}(\ell)},$ (9)
where
$\Delta\Lambda^{(q)}_{r,\ell}=\Lambda^{(q)}_{r,\ell}-\mu.$
[…]
The envelope relies on the transport factors $\Gamma^{(q)}_{t,\ell}$ obtained from the first-order diagonal expansion of the Jacobian product $\mathcal{M}_{t,\ell}$ (Appendix A). These factors admit a transparent structure: at zeroth order they reduce to products of gating values, which are the time scales operating in state space; the first-order terms add gate derivatives mixed with weight matrices (see Appendix B for architecture-specific derivations). The numerical analysis in [38] shows that truncating the matrix product to first order preserves numerical accuracy.
Alternatively, one could define the envelope directly from the full transport matrix as
$f_{\mathcal{M}}(\ell)=\sum_{q=1}^{H}|\Lambda^{(q)}[\mathcal{M}_{t,\ell}]_{qq}|,$
bypassing the first-order expansion. This would yield $H$ numbers with no structural decomposition into gating and mixing contributions, making it impossible to assign a time scale to individual neurons. The first-order expansion provides this interpretability without losing numerical quality: as verified in Appendix C, the envelope $f(\ell)$ in Eq. 8 preserves the decay profile of $f_{\mathcal{M}}(\ell)$ (Spearman $\rho\geq 0.979$, Pearson $r\geq 0.82$, mean $0.96$).
5.2 Gradient contribution at lag $\ell$
We begin by isolating the portion of the BPTT gradient that reflects the influence of past states. Throughout this section we fix a reference time $t$ and interpret $\ell\geq 1$ as a backward temporal displacement (lag) relative to $t$. All quantities indexed by $(t,\ell)$ refer to contributions originating from the state at lag $\ell$ prior to time $t$.
Let $h_{t}$ denote the recurrent hidden state used by BPTT (see Appendix B for architecture-specific details). Recalling the decomposition in Eq. (4), the lag-$\ell$ contribution to the parameter gradient can be written as
$g_{t,\ell}(\theta)\;=\;\delta_{t}^{\top}\mathcal{M}_{t,\ell}\,B_{\ell}(\theta)\;\in\;\mathbb{R}^{1\times P}.$ (10)
Then, decompose the parameter Jacobian into neuronwise rows,
$B_{\ell}(\theta)=\begin{bmatrix}B^{(1)}_{\ell}\\ \vdots\\ B^{(H)}_{\ell}\end{bmatrix},\qquad B^{(q)}_{\ell}\in\mathbb{R}^{1\times P},$
Following the first-order expansion of Jacobian products (see Appendix A for the general formulation and Appendix B for architecture-specific details), we approximate $\mathcal{M}_{t,\ell}$ to first order and retain only the diagonal components of the expansion. This yields neuronwise transport coefficients $\Gamma^{(q)}_{t,\ell}$, which quantify how gradient signals propagate through the recurrent dynamics. Combining these coefficients with the optimizer base rate (Section 4) produces the generalized effective learning rates $\mu^{(q)}_{t,\ell}$. Using the generalized effective learning rates $\mu^{(q)}_{t,\ell}$ yields the following first-order approximation of Eq. (10):
$g_{t,\ell}(\theta)\;\approx\;\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\delta_{t}^{(q)}\,B^{(q)}_{\ell}(\theta).$ (11)
As will be shown in the following subsections, this allows us to precisely determine the contribution of the envelope (8) in the detection analysis, which leads to the definition of the learnability window.
5.3 Binary detection and matched statistic
We formalize finite-horizon learnability by asking when information about a state at lag $\ell$ remains statistically present in the stochastic gradient at time $t$. The lag-$\ell$ contribution to the BPTT gradient is, up to first-order approximation, the vector $g_{t,\ell}(\theta)\in\mathbb{R}^{1\times P}$ defined in Eq. (11).
[…]
We therefore compress the vector-valued quantity $g_{t,\ell}(\theta)$ into a scalar by applying a fixed linear readout. Let $w\in\mathbb{R}^{P}$ be a random unit-norm vector, drawn independently of the data and fixed throughout training and across all model instances. This projection defines a one-dimensional observable that can be interpreted as a random but unbiased probe of the gradient geometry.
Applying this readout to $g_{t,\ell}(\theta)$ and using the first-order approximation Eq. (11), we obtain the scalar neuronwise alignment variables
$\zeta^{(q)}_{t,\ell}=\delta_{t}^{(q)}\,\langle B^{(q)}_{\ell}(\theta),\,w\rangle,$ (12)
so that the lag-$\ell$ contribution reduces to the weighted sum $\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\zeta^{(q)}_{t,\ell}$.
Because the same realization of $w$ is used for all models and training runs, differences in detectability across architectures reflect differences in temporal credit assignment and gradient transport, rather than artifacts of the projection itself. Moreover, since $w$ is drawn uniformly from the unit sphere, its distribution is rotationally invariant. As a result, the projection $\langle B^{(q)}_{\ell}(\theta),\,w\rangle$ provides an unbiased random probe of the typical magnitude and relative alignment of the neuronwise parameter-sensitivity directions, without privileging any fixed parameter direction. More generally, random projections preserve inner-product and norm structure in high dimension with high probability (the Johnson–Lindenstrauss phenomenon [27]), so the signal-to-noise ratio of the projected statistic faithfully reflects that of the full gradient.
The scalar variables $\zeta^{(q)}_{t,\ell}$ quantify how the transported gradient components associated with neuron $q$ align, through the parameter Jacobian, with the fixed readout direction $w$. The effective learning rates $\mu^{(q)}_{t,\ell}$ weight these alignment terms and thus govern how strongly each lag influences parameter updates.
Framed this way, the question of whether RNNs can use BPTT to exploit a dependency at lag $\ell$ becomes a question of detectability: is the signal induced by the state at lag $\ell$ large enough, relative to gradient noise, to remain statistically distinguishable? This viewpoint allows us to apply information-theoretic tools to characterize when past states remain recoverable during training.
Define the expected alignment $m_{q}(\ell)=\mathbb{E}[\zeta^{(q)}_{t,\ell}$] and construct a matched statistic that aggregates lag-$\ell$ evidence across neurons:
$S_{t,\ell}=\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,\mathrm{sgn}\!\big(m_{q}(\ell)\big)\,\zeta^{(q)}_{t,\ell}$. (13)
The factor $\mathrm{sgn}(m_{q}(\ell))$ flips each coordinate so that its contribution is oriented along the direction of its expected alignment. Consequently, under the data distribution, $\mathrm{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}$ has nonnegative mean, and every term in the sum contributes constructively. [^2] Formally, $\mathbb{E}\big[\mathrm{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}\big]=|m_{q}(\ell)|$ by definition of $m_{q}(\ell)$. Thus the expected signal contained in $S_{t,\ell}$ is always nonnegative, irrespective of the sign pattern of the raw alignments $m_{q}(\ell)$. The sign $\mathrm{sgn}(m_{q}(\ell))$ should be understood as a population-level orientation defining an ideal matched statistic; in practice, detectability depends only on the magnitude of the mean shift and is assessed from the empirical average of $S_{t,\ell}$, without requiring prior knowledge of this sign.
Taking expectations over the randomness of the data yields
$\mathbb{E}[S_{t,\ell}]=\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,|m_{q}(\ell)|=\overline{m}_{\mu}(\ell)\,f(\ell),$ (14)
where
$\overline{m}_{\mu}(\ell)=\frac{\sum_{q=1}^{H}\mu^{(q)}_{t,\ell}\,|m_{q}(\ell)|}{\sum_{q=1}^{H}\big|\mu^{(q)}_{t,\ell}\big|}$ (15)
is an envelope-normalized alignment coefficient aggregating the neuronwise magnitudes $|m_{q}(\ell)|$ with signed weights $\mu^{(q)}_{t,\ell}$.
This factorization separates two complementary aspects of temporal credit assignment: (i) $\overline{m}_{\mu}(\ell)$ captures the average informational alignment between the gradient signal and the parameter-sensitivity directions at lag $\ell$, and (ii) $f(\ell)$ quantifies the aggregate gain determining how strongly lagged gradient information contributes to parameter updates after both recurrent transport and optimizer scaling. Together, they determine the expected strength of the lagged contribution to the overall gradient, that is, how detectable past dependencies remain after being propagated through the coupled state–parameter dynamics of training.
Averaging $S_{t,\ell}$ over $N$ independent training sequences yields $\widehat{S}_{N}(\ell)=\tfrac{1}{N}\sum_{n=1}^{N}S^{(n)}_{t,\ell}$, the empirical matched statistic that forms the basis for the finite-sample analysis in the next section.
5.4 Finite-sample analysis
We derive the finite-sample requirements for detecting a lag-$\ell$ dependency from noisy gradient information.
5.4.1 Statistical model
[…]
We extend the zero-mean formulation of [49] to a two-parameter location family, allowing small shifts in the mean corresponding to the presence or absence of a detectable signal. At lag $\ell$, the averaged matched statistic $\widehat{S}_{N}(\ell)$ over $N$ independent training sequences is modeled as
$\widehat{S}_{N}(\ell)\;\big|\;B\;\sim\;\mathcal{S}\alpha\mathcal{S}\!\Big(\theta_{\text{out}},\;\sigma_{\alpha}(\ell)/N^{1-1/\alpha}\Big),\qquad\theta_{\text{out}}\!\in\!\left\{+\tfrac{1}{2}\Delta(\ell),\;-\tfrac{1}{2}\Delta(\ell)\right\},$ (16)
where $\theta_{\text{out}}$ is the location parameter, the mean separation $\Delta(\ell)=\overline{m}_{\mu}(\ell)f(\ell)$ quantifies the strength of the lag-$\ell$ signal, $\sigma_{\alpha}(\ell)$ acts as a noise-scale proxy, and $B\in\{0,1\}$ describes the non-detection/detection outcome.
For analytical tractability, we assume bounded alignment and noise scales, $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$ and $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$, which guarantee well-defined sample-complexity constants without affecting the asymptotic exponents. The factor $N^{1-1/\alpha}$ encodes the slow concentration typical of $\alpha$-stable averages, and a justification for this choice is discussed in Appendix E.
5.4.2 Information-theoretic bounds
[…]
$P_{\text{det}}=\mathcal{S}\alpha\mathcal{S}\!\Big({+}\tfrac{1}{2}\Delta(\ell),\,\sigma_{\alpha}(\ell)/N^{\,1-1/\alpha}\Big),\qquad P_{\text{non}}=\mathcal{S}\alpha\mathcal{S}\!\Big({-}\tfrac{1}{2}\Delta(\ell),\,\sigma_{\alpha}(\ell)/N^{\,1-1/\alpha}\Big),$
where $\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell)$ denotes the mean separation.
For $\alpha>1$, symmetric $\alpha$–stable location families satisfy local asymptotic normality (LAN) under suitable normalization, enabling the use of information-theoretic bounds [23, 34, 42]. In the present setting, the empirical statistic $\widehat{S}_{N}(\ell)$ defines a triangular array of location models [52] whose scale shrinks with the number of sequences as $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$. For such families, LAN holds with normalization rate $r_{N}=1/s_{N}$, which governs the detectability thresholds derived below. As a consequence, the log-likelihood ratio between $P_{\text{det}}$ and $P_{\text{non}}$ admits a quadratic expansion for small location shifts, yielding the lower bound for the Kullback-Leibler (KL) divergence
$D_{\mathrm{KL}}\!\left(P_{\text{det}}\|P_{\text{non}}\right)\;\geq\;c_{\alpha}\,\frac{N^{\,2(1-1/\alpha)}\,\Delta(\ell)^{2}}{\sigma_{\alpha}(\ell)^{2}},$ (17)
where $c_{\alpha}>0$ depends only on $\alpha$. A derivation of this bound for shrinking-scale $\alpha$-stable location models is provided in Appendix F.
Combining (17) with standard relations between KL divergence and mutual information [12] yields
$I\!\left(B;\widehat{S}_{N}(\ell)\right)\;\geq\;\min\!\left\{\log 2,\;c_{\alpha}\,\frac{N^{\,2(1-1/\alpha)}}{\sigma_{\alpha}(\ell)^{2}}\,\overline{m}_{\mu}(\ell)^{2}\,f(\ell)^{2}\right\}.$ (18)
Applying Fano’s inequality [12] to the binary detection problem yields the minimal number of independent sequences $N$ required to achieve a target error probability $P_{e}\leq\epsilon$:
$N\;\geq\;\left(\frac{\sigma_{\alpha}(\ell)}{\sqrt{c_{\alpha}}\,\overline{m}_{\mu}(\ell)f(\ell)}\right)^{\!\frac{\alpha}{\alpha-1}}\!\!\Big(\log\tfrac{1}{2\epsilon}\Big)^{\frac{\alpha}{2(\alpha-1)}}.$ (19)
This expression links sample complexity to the noise scale $\sigma_{\alpha}(\ell)$, the average alignment $\overline{m}_{\mu}(\ell)$, and the envelope $f(\ell)$. In contrast to the Gaussian case, the exponent $\alpha/(\alpha-1)>2$ for $\alpha<2$, reflecting the slower statistical concentration induced by heavy-tailed gradient noise. Full derivations of Eqs. (18) and (19) are given in Appendix G.
5.5 Learnability window and scaling laws
Throughout this section we adopt the corrected concentration exponent
$\kappa_{\alpha}\;=\;\frac{\alpha}{\alpha-1},\qquad\alpha>1,$ (20)
which governs the finite-sample detectability of lagged dependencies under $\alpha$-stable gradient noise when statistics are averaged over independent sequences. Equivalently, the noise floor of the empirical matched statistic decays as $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$.
Starting from the finite-sample condition in Eq. (19), we can invert the relation to express the minimal effective learning rate mass required for reliable detection at a given sample size $N$. Rearranging terms [^3] To obtain Eq. (21) from Eq. (19), one takes the $\kappa_{\alpha}$th root of the bound, isolates the factor $1/f(\ell)$, and inverts the (positive) inequality. This yields the minimal effective learning rate mass required for detectability at lag $\ell$. yields the per-lag detectability threshold
$\varepsilon_{\mathrm{th}}(\ell)=\frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)}\cdot\frac{1}{\sqrt{c_{\alpha}}}\,\sqrt{\log\!\Big(\tfrac{1}{2\epsilon}\Big)}.$ (21)
This threshold compactly summarizes the interplay of noise scale $\sigma_{\alpha}(\ell)$, data size $N$, and alignment strength $\overline{m}_{\mu}(\ell)$: for a dependency at lag $\ell$ to be statistically detectable under heavy-tailed $\alpha$-stable noise, the envelope $f(\ell)$ must exceed $\varepsilon_{\mathrm{th}}(\ell)$. In this formulation, smaller $\alpha$ values (heavier tails) increase $\kappa_{\alpha}$, thereby slowing statistical concentration to $N^{1/\kappa_{\alpha}}$ and reducing the effective temporal range over which informative gradients can be recovered.
We can now define the learnability window.
Definition 5.1 (Learnability window under $\alpha$-stable noise).
The learnability window is
$\mathcal{H}_{N}=\sup\big\{\ell\geq 1:\,f(\ell)\geq\varepsilon_{\mathrm{th}}(\ell)\big\}.$ (22)
Intuitively, $\mathcal{H}_{N}$ is the largest lag for which the transported gradient retains a recoverable signal. Even if Jacobians are numerically stable, once $f(\ell)$ falls below $\varepsilon_{\mathrm{th}}(\ell)$, heavy-tailed fluctuations dominate and credit assignment becomes statistically infeasible.
From the Fano bound in Eq. (19), the minimal number of independent sequences sufficient to detect a lag-$\ell$ dependency can be written as
$N(\ell)\;=\;\kappa_{\alpha,\epsilon}\,\left(\frac{\sigma_{\alpha}(\ell)}{\overline{m}_{\mu}(\ell)f(\ell)}\right)^{\!\kappa_{\alpha}},\qquad\kappa_{\alpha,\epsilon}\;=\;\frac{1}{c_{\alpha}^{\kappa_{\alpha}/2}}\,\Big(\log\tfrac{1}{2\epsilon}\Big)^{\kappa_{\alpha}/2}.$ (23)
The factor $\kappa_{\alpha,\epsilon}$ depends only on the tail index $\alpha$ and the target error level $\epsilon$ (fixed across lags) and will be absorbed into constant factors in our asymptotic statements.
Here, we formalize two regularities that drive the scaling of $\mathcal{H}_{N}$. Proofs appear in Appendix H.
Lemma 5.1 (Monotonicity in the lag).
Fix $t$ and a neuron $q$. Assume gate activations lie in $[0,1]$ and activation derivatives are bounded in $[0,1]$. Then the transport factors $\Gamma^{(q)}_{t,\ell}$ are nonincreasing in $\ell$, and consequently the envelope $f(\ell)$ is nonincreasing up to bounded optimizer scaling.
Lemma 5.2 (Sample-complexity scaling and window bounds).
Assume there exist constants $0<c_{\sigma}\leq C_{\sigma}$ and $0<c_{m}\leq C_{m}$ such that, over the lags of interest, $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$ and $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$. Then, with $N(\ell)$ defined in Eq. (23), there are constants $0<c_{\star}\leq C_{\star}$ for which
$c_{\star}\,f(\ell)^{-\kappa_{\alpha}}\ \leq\ N(\ell)\ \leq\ C_{\star}\,f(\ell)^{-\kappa_{\alpha}},\qquad c_{\star}=\kappa_{\alpha,\epsilon}\Big(\tfrac{c_{\sigma}}{C_{m}}\Big)^{\!\kappa_{\alpha}},\quad C_{\star}=\kappa_{\alpha,\epsilon}\Big(\tfrac{C_{\sigma}}{c_{m}}\Big)^{\!\kappa_{\alpha}}.$ (24)
Consequently, with $f^{\leftarrow}(y)=\sup\{\ell\geq 1:\ f(\ell)\geq y\}$ being the generalized inverse of $f(\ell)$,
$f^{\leftarrow}\!\Big(\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}\Big)\ \leq\ \mathcal{H}_{N}\ \leq\ f^{\leftarrow}\!\Big(\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\Big).$ (25)
The first lemma captures the monotone, multiplicative attenuation induced by gates; the second makes explicit that the per-lag sample complexity grows as the inverse $\kappa_{\alpha}$-power of the envelope $f(\ell)$. Together, they reduce the problem of characterizing $\mathcal{H}_{N}$ to inverting the decay profile of $f(\ell)$ at the level $N^{-1/\kappa_{\alpha}}$. In other words, once $\sigma_{\alpha}(\ell)$ and $\overline{m}_{\mu}(\ell)$ are bounded, the temporal reach of learnability is controlled by the speed with which $f(\ell)$ decays with $\ell$ and by the heavy-tailed concentration exponent $\kappa_{\alpha}$.
Canonical envelope scaling regimes.
In practice, gated RNNs may exhibit intricate envelope profiles, including mixtures of decay regimes or multi-phase behavior across different lag ranges. To develop analytic insight, we focus on three canonical asymptotic scaling profiles: logarithmic, polynomial, and exponential decay. These regimes provide a structural classification of how $f(\ell)$ can attenuate in the large-lag limit and form the basis of the corresponding learnability scaling laws. Real networks need not follow a single regime uniformly across all lags; finite-dimensional implementations and bounded gates inevitably induce crossover behavior, typically resulting in exponential-like cutoffs at sufficiently large $\ell$. More complex envelope shapes can nevertheless be analyzed within the same framework by decomposing them into these canonical asymptotic components.
For each of these canonical asymptotic regimes of $f(\ell)$, we obtain the corresponding large-$\ell$ scaling laws (up to multiplicative constants):
  1. (i)
    Logarithmic decay: if $f(\ell)\!\asymp\!c/\log(1+\ell)$, then $N(\ell)\!\asymp\![\log(1+\ell)]^{\kappa_{\alpha}}$ and $\mathcal{H}_{N}\!\asymp\!\exp(\kappa N^{1/\kappa_{\alpha}})-1$.
  2. (ii)
    Polynomial decay: if $f(\ell)\!\asymp\!c\,\ell^{-\beta}$, then $N(\ell)\!\asymp\!\ell^{\kappa_{\alpha}\beta}$ and $\mathcal{H}_{N}\!\asymp\!N^{1/(\kappa_{\alpha}\beta)}$.
  3. (iii)
    Exponential decay: if $f(\ell)\!\asymp\!c\,\lambda^{\ell}$ with $\lambda\!\in(0,1)$, then $N(\ell)\!\asymp\!\lambda^{-\kappa_{\alpha}\ell}$ and $\mathcal{H}_{N}\!\asymp\!(\log N)/[\kappa_{\alpha}\log(1/\lambda)]$.
In these expressions, the symbol $\asymp$ denotes equality at the level of asymptotic order. These scaling laws characterize how gating geometry and heavy-tailed noise jointly determine the large-lag behavior of temporal learnability. Exponential decay corresponds to rapid forgetting and only logarithmic growth of the horizon; polynomial decay yields algebraic horizon growth; and logarithmic decay represents a boundary regime with rapidly expanding $\mathcal{H}_{N}$. In all cases, smaller $\alpha$ compresses $\mathcal{H}_{N}$ by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Full derivations and constant factors are provided in Appendix I.
5.6 Theoretical implications
Combining the Fano-derived sample-complexity bound in Eq. (19) with the learnability window analysis of Sec. 5.5 yields several theoretical insights into how gating and optimizer structure, noise statistics, and finite data interact to determine temporal learnability.
(i) Envelopes set sample complexity under heavy-tailed noise.
The detectability of a lag-$\ell$ dependency is governed by the signal amplitude $\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell)$, which grows linearly with the envelope $f(\ell)$ (see Eq. (14)). In the mutual information bound of Eq. (18), this signal enters quadratically as $\Delta(\ell)^{2}$, while the contribution of data enters through the concentration factor $N^{2(1-1/\alpha)}=N^{2/\kappa_{\alpha}}$.
When $\alpha<2$, the dependence on $N$ is sublinear, reflecting the slow statistical concentration of heavy-tailed averages. As a result, at fixed lag $\ell$ and up to bounded prefactors associated with $\overline{m}_{\mu}(\ell)$ and $\sigma_{\alpha}(\ell)$, slower envelope decay yields a multiplicative gain in signal strength that dominates the effect of increasing the dataset size. This imbalance is made explicit by the Fano bound (19), which yields the master proportionality
$N(\ell)\;\propto\;f(\ell)^{-\kappa_{\alpha}}.$ (26)
Equation (26) identifies the envelope $f(\ell)$ as the primary control variable governing sample complexity under heavy-tailed noise. Concrete scaling laws follow once this proportionality is combined with a specific decay regime for $f(\ell)$. For example, polynomial envelope decay yields algebraic growth of $N(\ell)$ and rapidly expanding learnability windows, whereas exponential envelope decay leads to rapidly increasing sample requirements and slowly-increasing horizons; see Appendix I for details.
(ii) The tail index $\alpha$ governs statistical efficiency.
The concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$ quantifies how rapidly noise averages out when statistics are aggregated across independent sequences. For Gaussian fluctuations ($\alpha=2$), one recovers the familiar concentration rate $N^{-1/2}$; for $\alpha<2$, the slower rate $N^{-1/\kappa_{\alpha}}=N^{-(1-1/\alpha)}$ substantially weakens the effective signal-to-noise ratio. Thus, even for identical gating geometry and identical envelopes $f(\ell)$, smaller values of $\alpha$ increase $\kappa_{\alpha}$, compress the learnability window, and raise the number of samples required for reliable detection.
Importantly, this dependence implies that the effective scaling regime is determined jointly by the envelope $f(\ell)$ and the noise statistics, rather than by architectural structure alone.
(iii) Architectural and training implications.
Architectures that permit slowly decaying envelopes are intrinsically capable of supporting long-range learnability under heavy-tailed noise. At the same time, such behavior is not guaranteed by architecture alone: training dynamics and noise statistics determine which decay regime is effectively realized. Conversely, training methods that reduce the impact of heavy-tailed fluctuations, such as clipping, normalization, or stabilizing adaptive optimizers, effectively increase $\alpha$, thereby reducing $\kappa_{\alpha}$ and improving statistical efficiency without altering model capacity.
(iv) Vanishing learnability as an information-theoretic obstruction.
The learnability window $\mathcal{H}_{N}$ may vanish identically for certain architectures, meaning that $\mathcal{H}_{N}=0$ for all finite $N$. In this case, the envelope $f(\ell)$ falls below the statistical detectability threshold for every $\ell\geq 1$, and the corresponding sample complexity $N(\ell)$ diverges. This does not reflect an optimization failure or insufficient data, but an information-theoretic obstruction: the transported gradients carry no statistically usable signal at nonzero lags. Equivalently, the KL divergence between the detection and non-detection models vanishes, so no estimator can reliably distinguish the two hypotheses regardless of sample size.
6 Empirical validation
The purpose of this section is not to compare architectures in terms of predictive performance, but to empirically probe the mechanisms identified by the theory: how envelope decay, noise statistics, gating structure, and optimizer adaptation jointly determine finite-horizon learnability.
Here, learnability refers to the statistical detectability of lagged gradient contributions during training. Detectability is a necessary (though not sufficient) condition for successful learning of long-range dependencies: if the gradient signal originating from a past state is not statistically distinguishable from noise, consistent credit assignment cannot occur even when gradients remain numerically stable.
Accordingly, the experiments are designed to test the structural predictions of the theory. In particular, we aim to: (i) measure the empirical decay profiles of the envelope $f(\ell)$ and assess how they align with the canonical decay regimes considered by the theory; (ii) quantify how the resulting empirical learnability windows $\widehat{\mathcal{H}}_{N}$ vary with gating structure, optimizer dynamics, and sample size, and examine whether their scaling behavior is consistent with the regime implied by the observed envelope geometry; and (iii) investigate how the neuronwise time-scale spectra $\{\tau_{q}\}_{q}$ and the estimated gradient-noise statistics jointly relate to the empirically realized learnability regimes.
6.1 Experimental setup
Task and data.
We consider a synthetic regression task defined on input–output pairs $\{(x_{t},y_{t})\}_{t=1}^{T}$, where $x_{t}\in\mathbb{R}^{16}$ and the target $y_{t}\in\mathbb{R}$ is scalar. All experiments use sequences of length $T=1024$.
The target is generated as a weighted sum of delayed inputs,
$y_{t}=\sum_{k=1}^{5}c_{k}\,u^{\top}x_{t-\ell_{k}}+\varepsilon_{t},$ (27)
where $u\in\mathbb{R}^{16}$ is a random unit vector, the task lags are $\{\ell_{k}\}=\{32,\,64,\,128,\,192,\,256\}$, the mixing coefficients are $\{c_{k}\}=\{0.6,\,0.5,\,0.4,\,0.32,\,0.26\}$, and $\varepsilon_{t}$ is independent Gaussian noise with standard deviation $\sigma_{\mathrm{noise}}=0.3$. These delayed components induce informative dependencies spanning short, intermediate, and relatively long temporal scales.
Each model is trained on $8000$ independent sequences. All effective learning-rate and envelope statistics are computed on an independent diagnostic set of $8000$ fresh sequences generated from the same data-generating process. No sequences are reused across training and diagnostic splits.
To assess temporal credit assignment independently of the specific task lags, all diagnostic quantities are evaluated on a separate lag grid $\ell\in[4,256]$ consisting of $128$ uniformly spaced values. This grid covers both sub-task and supra-task temporal scales, allowing us to probe the full decay profile of the learned envelope.
Architectures.
We consider five recurrent architectures described in Appendix B, spanning increasing gating expressivity and potential time-scale heterogeneity: ConstGate, SharedGate, DiagGate, GRU, and LSTM. All models use hidden dimension $H=128$ and a linear readout $y_{t}=w^{\top}h_{t}$ trained with mean-squared error loss. Recurrent weight matrices are initialized orthogonally, biases are initialized to zero unless otherwise stated, and gate pre-activations in learned-gate models are initialized near zero, yielding initial gate values close to $\tfrac{1}{2}$ and well-conditioned initial time scales, consistent with the time-scale initialization heuristic of [38].
For brevity, in figures and legends we denote ConstGate, SharedGate, DiagGate, LSTM, and GRU by const, shared, diag, lstm, and gru, respectively.
Training protocol.
All models are trained using mini-batch AdamW [39] with base learning rate $\mu=10^{-3}$, weight decay $\lambda=10^{-4}$, batch size $B=512$, and gradient clipping at $\ell_{2}$ norm $1.0$. Training uses $8000$ independent sequences of length $T=1024$ for $500$ epochs. No learning-rate schedule is applied, and all architectures share identical optimization hyperparameters. Adaptive optimizers such as AdamW are known to preserve heavy-tailed gradient fluctuations even in the presence of gradient clipping [8]. We verified that disabling gradient clipping yields qualitatively similar envelope and noise statistics (results not shown), indicating that clipping affects numerical stability but not the scaling regimes of interest.
To clarify the role of the optimizer in shaping envelope geometry and the learnability window, additional experiments with plain SGD and SGD with momentum are reported in Appendix J.
Diagnostic evaluation.
After training, parameters are frozen and all diagnostics are computed on an independent set of $8000$ fresh sequences generated from the same data-generating process, with no overlap in sequences or random seeds. For each trained model, the transport factors $\Gamma^{(q)}_{t,\ell}$ are computed from the frozen gate activations using the architecture-specific expressions derived in Appendix B. For AdamW, the adaptive base rates $\Lambda^{(q)}_{r,\ell}$ are obtained from the Rayleigh projection (7); for SGD they reduce to the constant $\Lambda^{(q)}_{r,\ell}=\mu$. For each diagnostic lag $\ell$ and unit $q$, the magnitudes $|\mu_{t,\ell}^{(q)}|$ are averaged over time steps and independent sequences. The empirical envelope is then defined as
$\hat{f}(\ell)=\sum_{q}\Big\langle|\mu_{t,\ell}^{(q)}|\Big\rangle,$
which corresponds to the empirical $\ell_{1}$ aggregation of neuronwise effective learning rates at lag $\ell$.
Heavy-tailed statistics.
[…]
Empirical learnability window.
The Fano-based detectability threshold (21) can be rewritten as a signal-to-noise ratio (SNR) criterion. Substituting $\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell)$ and rearranging, the condition $f(\ell)\geq\varepsilon_{\mathrm{th}}(\ell)$ in the learnability window (22) is equivalent to
$\mathrm{SNR}(\ell,N)\;:=\;\frac{|\Delta(\ell)|\;N^{1-1/\alpha}}{\sigma_{\alpha}(\ell)}\;\geq\;\underbrace{\frac{1}{\sqrt{c_{\alpha}}}\,\sqrt{\log\!\bigl(\tfrac{1}{2\epsilon}\bigr)}}_{\displaystyle=:\,\varepsilon},$ (28)
where the right-hand side is a positive constant depending only on $\alpha$ and the error probability $\epsilon$. This reformulation makes explicit that the Fano bound constrains a single scalar quantity, i.e. the SNR, and that the (aggregated) detection threshold $\varepsilon$ affects only the absolute position of $\mathcal{H}_{N}$, not how it scales with $N$.
In practice, we estimate the SNR by substituting the empirical signal $|\widehat{\overline{m}}_{\mu}(\ell)|$, the estimated tail index $\hat{\alpha}$, and the estimated scale $\hat{\sigma}_{\alpha}(\ell)$:
$\widehat{\mathrm{SNR}}(\ell,N)\;=\;\frac{|\widehat{\overline{m}}_{\mu}(\ell)|\;N^{1-1/\hat{\alpha}}}{\hat{\sigma}_{\alpha}(\ell)}.$
Since $c_{\alpha}$ is an existence constant from the LAN expansion (Appendix F) and is not available in closed form, we set the detection threshold $\varepsilon$ directly. This preserves the full scaling structure of the theoretical bound (the concentration rate $N^{1-1/\alpha}$, the dependence on signal amplitude and noise scale) while replacing the single unknown constant with a user-specified parameter. The empirical learnability window is then
$\widehat{\mathcal{H}}_{N}\;=\;\max\bigl\{\ell:\widehat{\mathrm{SNR}}(\ell,N)>\varepsilon\bigr\}.$ (29)
Time-scale extraction.
The effective learning rates $\mu_{t,\ell}^{(q)}$ attenuate the gradient contribution of neuron $q$ as a function of lag. To summarize each neuron’s temporal reach with a single scalar, we fit the exponential model
$|\mu_{t,\ell}^{(q)}|\;\approx\;C_{q}\exp(-\ell/\tau_{q})$
[…]
6.2 Results
Envelope scaling.
[…]
ConstGate and SharedGate display rapid envelope decay. On a semi-logarithmic scale (middle), both curves are well approximated by straight lines, consistent with exponential attenuation. Exponential fits of the form $\hat{f}(\ell)\approx c\,\exp(-\lambda\ell)$ yield decay rates $\lambda_{\text{const}}\approx 0.68$ and $\lambda_{\text{shared}}\approx 0.39$, with coefficients of determination $r^{2}\approx 0.99$ and $r^{2}\approx 0.99$, respectively. Power-law fits perform substantially worse ($r^{2}\approx 0.86$ and $r^{2}\approx 0.85$), confirming that these architectures lie firmly in the exponential regime. Their envelopes decay by several orders of magnitude within a few tens of lags, indicating a short effective temporal range.
DiagGate exhibits qualitatively slower attenuation. An exponential fit yields a small decay rate $\lambda_{\text{diag}}\approx 0.020$ ($r^{2}\approx 0.99$), corresponding to a very long effective time scale. However, on a log–log scale (right), the envelope displays an approximately linear trend over the diagnostic range. A power-law fit $\hat{f}(\ell)\approx c\,\ell^{-\beta}$ yields $\beta_{\text{diag}}\approx 1.71$ with $r^{2}\approx 0.93$. The exponential and algebraic models therefore provide comparable descriptions over the finite lag window explored in our experiments. In DiagGate, the exponential fit is characterized by an extremely small decay rate, so that over the observed range of lags the curve is nearly indistinguishable from a power-law trend. We therefore interpret DiagGate as operating in an intermediate regime: empirically consistent with approximate polynomial scaling across the explored window, yet theoretically compatible with an eventual exponential cut-off at sufficiently large lags.
GRU and LSTM extend this heterogeneous behavior. For GRU, exponential and power-law fits achieve comparable goodness of fit ($\lambda_{\text{gru}}\approx 0.018$, $r^{2}\approx 0.97$; $\beta_{\text{gru}}\approx 1.59$, $r^{2}\approx 0.96$), indicating a very slow exponential attenuation. Over the finite lag window explored, this shallow exponential is empirically difficult to distinguish from a power-law trend, although the two imply different asymptotic scaling regimes. For LSTM, the distinction is sharper: the exponential fit is comparatively weaker ($\lambda_{\text{lstm}}\approx 0.011$, $r^{2}\approx 0.91$), whereas the power-law fit yields $\beta_{\text{lstm}}\approx 1.01$ with $r^{2}\approx 0.99$. Over the diagnostic range, the LSTM envelope therefore aligns closely with polynomial scaling, with an exponential cut-off necessarily emerging at sufficiently large lags due to finite-state dynamics.
The interpretation of an intermediate algebraic-like regime for DiagGate, GRU, and LSTM is further supported by the decay of the complementary cumulative distribution function (CCDF) of the neuronwise effective time scales, reported below. In these architectures, the tail of the CCDF exhibits a slow decay compatible with a broad distribution of time constants, followed by a clear attenuation at large scales. This pattern is consistent with an approximate power-law–like behavior over intermediate lags, combined with an inevitable exponential cut-off induced by finite-dimensional state dynamics.
Collectively, these results reveal a hierarchy of temporal regimes: homogeneous scalar gating produces fast exponential decay; diagonal gating induces slow, approximately algebraic attenuation; and multi-gate architectures such as GRU and LSTM realize broad mixtures of time scales that manifest as approximate power-law scaling over intermediate horizons. Importantly, the corresponding state Jacobians remain numerically well-conditioned in all cases (results not shown), indicating that Jacobian stability alone does not determine finite-horizon learnability. Rather, it is the geometry of the envelope that governs how strongly lagged gradient contributions influence parameter updates across temporal lags. All models exhibit consistent training loss reduction and stable optimization trajectories (learning curves not shown), confirming that they successfully fit the task under the shared training protocol. However, the objective of this study is not to fine-tune architectures for maximal predictive accuracy, but to characterize the structural conditions under which long-range dependencies remain statistically detectable.
(a) Linear scale
(b) Semi-log scale
(c) Log-log scale
[…]
Empirical learnability windows.
[…]
DiagGate displays qualitatively different behavior. For small $N$, the learnability window is negligible. Once the budget exceeds a few hundred sequences, $\widehat{\mathcal{H}}_{N}$ expands in discrete steps, progressively reaching intermediate and longer lags and stabilizing near $\widehat{\mathcal{H}}_{N}\approx 120$ for the largest budgets considered. This data-dependent expansion is consistent with the observed approximately algebraic envelope regime: slow attenuation allows additional lags to cross the detectability threshold as $N$ increases.
GRU and LSTM exhibit an even more pronounced extension of the horizon. Both architectures display a sharp transition: below a critical sample size the horizon is effectively zero, whereas above this threshold the window rapidly expands to the maximal diagnostic lag ($\widehat{\mathcal{H}}_{N}\approx 256$). This behavior reflects their broad mixture of time scales and slow envelope decay over intermediate lags. In these models, once the training budget exceeds the noise-dominated regime, a large portion of the temporal range becomes detectable within the explored horizon.
Taken together, these results confirm the structural prediction of the theory: exponential envelope decay leads to early saturation of the learnability window, whereas slow, approximately algebraic attenuation enables systematic expansion of the horizon with increasing data. These behaviors are precisely those implied by the master proportionality $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, which links envelope geometry to sample complexity under heavy-tailed noise.
[…]
Time-scale spectra.
To connect the envelope geometry back to gating, we inspect the distribution of neuronwise time scales $\{\tau_{q}\}_{q}$ inferred from the effective learning rates. Rather than relying solely on density estimates, we summarize these spectra via the CCDF, $\mathbb{P}(\tau_{q}\geq\tau)$, which directly quantifies the fraction of units operating on time scales longer than $\tau$. Figure 3 reports the empirical CCDFs for all architectures on a log-scaled $\tau$ axis, which makes heterogeneity in the slow tail immediately visible.
ConstGate exhibits an essentially degenerate spectrum, with all units clustered around a single characteristic time scale. SharedGate displays a similarly narrow spectrum, with only minor unit-to-unit variability. In contrast, DiagGate produces a broad and heterogeneous time-scale distribution: most units operate on short time scales, while a non-negligible fraction populates intermediate ranges and a small subpopulation extends to substantially longer $\tau$ values. GRU and LSTM further accentuate this heterogeneity, exhibiting particularly slowly decaying CCDF tails over intermediate $\tau$ together with a clear attenuation at the largest time scales. This mixed pattern, slow tail decay over an intermediate range followed by an effective cut-off, is consistent with the envelope behavior reported above: architectures with broader spectra and a slow tail support substantially extended temporal transport over finite horizons, whereas narrowly concentrated spectra correlate with rapid exponential envelope decay and early saturation of $\widehat{\mathcal{H}}_{N}$. Additional tail-emphasis visualizations and per-model density diagnostics are reported in Appendix K, which further illustrate the heterogeneous multi-scale structure underlying these regimes.
Figure 3: Complementary cumulative distribution functions (CCDFs) of the neuronwise effective time scales $\tau_{q}$. ConstGate and SharedGate exhibit narrowly concentrated spectra, whereas DiagGate, GRU, and LSTM display broad heterogeneous distributions with a slowly decaying intermediate tail and a visible attenuation at large $\tau$, consistent with a mixed multi-scale regime.
Noise statistics and selection of scaling regimes.
[…]
Importantly, these results indicate that the realized scaling behavior is not determined by architecture alone. Rather, gating geometry defines the space of accessible time-scale spectra, while the interaction between training dynamics and emergent noise statistics selects which scaling regime is stabilized in practice.
[…]
Figure 5: Estimated lag-dependent noise scale $\hat{\sigma}_{\alpha}(\ell)$ for all architectures (log scale). ConstGate and SharedGate exhibit rapidly decaying noise scales, indicating fast statistical concentration of gradient fluctuations. By contrast, DiagGate, GRU, and LSTM maintain substantially larger fluctuation levels across lags, with markedly slower decay. Such persistent noise at long temporal distances increases the statistical cost of rapidly vanishing envelopes and is consistent with the emergence of slower, more robust scaling regimes.
7 Discussion and future directions
This work develops a theoretical framework linking gating geometry, optimizer adaptation, heavy-tailed gradient noise, and the finite-horizon learnability of recurrent neural networks. The central message is that the generalized effective learning rates, rather than Jacobian stability in isolation, determine how much gradient signal remains statistically detectable across temporal lags. The decay profile of the envelope, together with the tail index $\alpha$, governs both sample-complexity requirements and the learnability window.
Our empirical results support this theoretical picture while clarifying the roles of architectural inductive bias and optimizer adaptation. Architectures with constrained gating tend to induce narrow and nearly synchronized time-scale spectra, leading to fast exponential envelope decay and short, largely data-insensitive learnability windows. Architectures capable of neuronwise or multi-gate modulation admit heterogeneous mixtures of time scales and can realize slower attenuation regimes, within which the learnability window expands systematically with available data. Adaptive optimizers amplify this differentiation: neurons with slow transport accumulate larger second-moment estimates, lowering their adaptive base rates, while fast neurons receive higher rates. This partial equalization of effective rates across neurons with different intrinsic time scales is absent under SGD. The decisive factor is therefore not nominal architectural complexity, but the dispersion of effective time scales that emerges from the joint dynamics of gating and optimization.
The emergence of algebraic-like envelope behavior should be interpreted as a finite operational regime rather than as an asymptotic property. In finite-dimensional recurrent systems, true scale invariance over arbitrarily large lags is structurally impossible. Even when a broad hierarchy of effective time scales induces approximately polynomial decay over an extended intermediate range, finite-size effects ultimately impose a crossover once the signal falls below the detectability threshold. At the same time, the realized scaling regime is not deterministic. Even architectures with rich gating structure may converge to synchronized time scales and revert to the exponential regime, depending on training dynamics, optimization hyperparameters, and task statistics. The architecture–optimizer pair defines the space of accessible time-scale spectra, but does not uniquely determine which regime is selected. Within this classification, the logarithmic decay regime represents a formal boundary case separating algebraic from exponential forgetting. We conjecture that it cannot be stably maintained in finite recurrent networks without substantial fine-tuning of the time-scale spectrum.
Heavy-tailed noise as a viability constraint.
The scaling laws make precise the statistical cost of each decay regime: exponential envelope decay entails exponential sample complexity, while polynomial decay entails only polynomial sample complexity. For any finite training budget, exponential forgetting rapidly exhausts the available statistical resources, whereas polynomial forgetting keeps long-range dependencies within reach. Heavy-tailed gradient fluctuations exacerbate this disparity by increasing $\kappa_{\alpha}$ and slowing statistical concentration. Heavy-tailed noise therefore acts as a viability constraint, disfavoring rapid forgetting and effectively selecting, among the dynamical regimes accessible to an architecture–optimizer pair, those that preserve signal over longer temporal scales.
We conjecture that this viability constraint induces a form of constrained self-organization: during training, recurrent networks—and potentially deep learning models more broadly—tend to evolve toward regimes of slower envelope decay in order to maintain learnability under their own gradient-noise statistics. This does not require explicit regularization; it follows from the fact that configurations with exponential forgetting become statistically unlearnable at long lags, creating implicit selection pressure toward slower decay. The achievable regime depends on the architecture–optimizer pair: constrained gating limits the accessible spectra, while flexible gating combined with adaptive optimization permits broader time-scale mixtures. This conjecture is consistent with our empirical observation that slow-decay regimes systematically co-occur with heavier-tailed gradient noise, whereas synchronized exponential regimes operate closer to Gaussian concentration.
Broader implications.
These observations highlight three broader implications. First, the temporal horizon of learnability is primarily controlled by the joint action of the envelope decay and optimizer adaptation. Second, the architecture–optimizer pair shapes the space of attainable time-scale spectra, biasing—but not guaranteeing—the emergence of heterogeneous dynamics that support long-range credit assignment. Third, under heavy-tailed gradient noise, increasing the number of training sequences yields diminishing returns for configurations operating in the exponential regime, whereas modifying the architecture–optimizer pair can alter the scaling class of temporal learnability entirely.
More broadly, the framework is not tied to a specific architecture or optimizer. We focused on gated recurrent networks because their dynamics admit a transparent interpretation in terms of tunable time scales, but similar limitations are expected in any learning system where signals propagate through long chains of Jacobians, including deep or sequence models where temporal depth can be traded for layer depth.
Finally, the present framework establishes which scaling regimes are statistically viable under finite data and heavy-tailed noise, but does not explain how training dynamics navigate the space of accessible regimes. A natural next step is the development of a generative stochastic model describing the dynamics of the effective learning rates during training—a theory of how time-scale spectra emerge, broaden, or collapse as the architecture–optimizer system co-evolves. Such a dynamical theory would complement the static viability analysis presented here with an account of long-range memory formation, addressing the open question of why some trained networks self-organize into broad time-scale mixtures while others collapse to synchronized dynamics.
Appendix A Matrix product expansion via the Fréchet derivative formulation
This section summarizes the first-order expansion of a product of matrices with structured perturbations introduced in [38].
Definition A.1 (Fréchet differentiability [32, 19]).
Let $f:\mathbb{C}^{n\times n}\to\mathbb{C}^{n\times n}$. We say that $f$ is Fréchet differentiable at $A\in\mathbb{C}^{n\times n}$ if there exists a bounded linear mapping $L_{f}(A,\cdot)$ such that
$\lim_{\|E\|\to 0}\frac{\|\,f(A+E)-f(A)-L_{f}(A,E)\,\|}{\|E\|}=0.$ (30)
If $f$ is Fréchet differentiable at $A$,
$f(A+E)\;=\;f(A)\;+\;L_{f}(A,E)\;+\;o(\|E\|).$ (31)
If $g$ and $h$ are Fréchet differentiable at $A$ and $f(X)=g(X)h(X)$, then
$L_{gh}(A,E)\;=\;L_{g}(A,E)\,h(A)\;+\;g(A)\,L_{h}(A,E).$ (32)
Consider the perturbed matrix product
$F(\varepsilon)=\prod_{j=1}^{n}\big(A_{j}+\varepsilon B_{j}\big),$ (33)
where $A_{j},B_{j}\in\mathbb{C}^{d\times d}$. By recursive application of (32) one obtains
$L_{F_{n}}(0,E)\;=\;\sum_{i=1}^{n}\left(\prod_{j=1}^{i-1}A_{j}\right)\,B_{i}\,\left(\prod_{j=i+1}^{n}A_{j}\right),$ (34)
and the first-order expansion
$F(\varepsilon)\;=\;\left(\prod_{j=1}^{n}A_{j}\right)\;+\;\varepsilon\sum_{m=1}^{n}\left(\prod_{j=1}^{m-1}A_{j}\right)\,B_{m}\,\left(\prod_{j=m+1}^{n}A_{j}\right)\;+\;O(\varepsilon^{2}).$ (35)
Ordering in BPTT.
The expansion is applied to Jacobian products arising in BPTT, where factors are naturally ordered backwards in time. In that setting, the indices are reparameterized so that the rightmost factor corresponds to the earliest time step. The algebraic structure of the first-order expansion remains identical, but the products appear with reversed temporal indexing.
Scope and limitations.
The diagonal first-order expansion defines the per-neuron effective learning rates used throughout the main text. The numerical accuracy of this expansion as an approximation of the full Jacobian product, including truncation error analysis and simulations across architectures, is validated in the original paper [38], which also discusses its limitations. Appendix C provides complementary evidence that the resulting envelope preserves the decay profile required by the learnability theory.
Appendix B RNNs: One-step Jacobians and effective learning rates
This appendix provides the full architecture-specific derivations of one-step Jacobians and per-neuron effective learning rates for the recurrent models considered in this paper. The derivations follow from the first-order expansion of Jacobian products summarized in Appendix A. For each architecture, we decompose the one-step Jacobian into a zeroth-order transport operator and a recurrent mixing correction, apply the first-order product rule, and extract the diagonal components that define the neuron-wise transport factors.
The effective learning rates are presented here for plain SGD with a fixed global step size $\mu$, following the original formulation in [38]. This keeps the derivations focused on the recurrent transport mechanisms that generate the neuron-wise factors $\Gamma^{(q)}_{t,\ell}$. The generalized formulation used in the main text, which replaces the global step size $\mu$ by the neuron-specific base rate $\Lambda^{(q)}_{r,\ell}$, is introduced in Section 4.
Section B.1 fixes notation. Sections B.2 and B.3 derive the effective learning rates for the LSTM and GRU, respectively. Section B.4 treats the baseline gated RNN architectures (DiagGate, SharedGate, ConstGate). Section B.5 concludes with a discussion of how these effective learning rates reveal an implicit multi-rate optimizer structure induced by gating.
B.1 Notation and conventions
Let $x_{t}\in\mathbb{R}^{D}$ be the input at step $t$, $h_{t}\in\mathbb{R}^{H}$ the hidden state, and, for the LSTM, $c_{t}\in\mathbb{R}^{H}$ the cell state. For any $v\in\mathbb{R}^{H}$, let $D(v)=\mathrm{diag}(v)\in\mathbb{R}^{H\times H}$. Hadamard (elementwise) product is denoted by $\odot$. The logistic and hyperbolic tangent functions are applied elementwise: $\sigma(\cdot)$ and $\tanh(\cdot)$. We use the diagonal slope matrices
$S^{\sigma}(u)=D(\sigma^{\prime}(u)),\qquad S^{\tanh}(u)=D\!\bigl(1-\tanh^{2}(u)\bigr).$
When convenient, $D(h_{t-1})$ is abbreviated by $D_{h,t-1}$. Weight matrices have shapes $W_{\bullet}\in\mathbb{R}^{H\times D}$, $U_{\bullet}\in\mathbb{R}^{H\times H}$, and $b_{\bullet}\in\mathbb{R}^{H}$.
Throughout, one-step Jacobians are derivatives with respect to the previous recurrent state. For the LSTM we stack $s_{t}=[h_{t};c_{t}]\in\mathbb{R}^{2H}$ and write
$J_{t}=\frac{\partial s_{t}}{\partial s_{t-1}}\in\mathbb{R}^{2H\times 2H},$
whereas for the GRU and the baseline gated RNNs we write
$J_{t}=\frac{\partial h_{t}}{\partial h_{t-1}}\in\mathbb{R}^{H\times H}.$
For a square matrix $A\in\mathbb{R}^{n\times n}$, $\mathrm{diagvec}(A)\in\mathbb{R}^{n}$ denotes the vector formed by the diagonal of $A$. We use interval-product notation throughout: for any matrix-valued sequence $\{A_{j}\}$,
$A_{a:b}=\prod_{j=b+1}^{a}A_{j},\qquad A_{b:b}=I.$
In the BPTT setting, this convention is applied to Jacobian products in the form
$\mathcal{M}_{t,\ell}=\prod_{j=\ell+1}^{t}J_{j},$
so that the rightmost factor corresponds to the earliest time step, consistent with the ordering convention stated in Appendix A.
In the LSTM case, when $\mathcal{M}_{t,\ell}$ is written in $2H\times 2H$ block form for $s=[h;c]$, the notation $\cdot\,]_{h,c}\in\mathbb{R}^{H\times H}$ refers to the top-right block mapping $c_{\ell}\mapsto h_{t}$.
B.2 LSTM: one-step Jacobian and effective learning rates
Dynamics.
An LSTM maintains $(h_{t},c_{t})$ and computes
$a^{i}_{t}$ $=W_{i}x_{t}+U_{i}h_{t-1}+b_{i},$ $i_{t}$ $=\sigma(a^{i}_{t}),$ (input gate) (36)
$a^{f}_{t}$ $=W_{f}x_{t}+U_{f}h_{t-1}+b_{f},$ $f_{t}$ $=\sigma(a^{f}_{t}),$ (forget/retention) (37)
$a^{o}_{t}$ $=W_{o}x_{t}+U_{o}h_{t-1}+b_{o},$ $o_{t}$ $=\sigma(a^{o}_{t}),$ (output/expression) (38)
$a^{g}_{t}$ $=W_{g}x_{t}+U_{g}h_{t-1}+b_{g},$ $g_{t}$ $=\tanh(a^{g}_{t}),$ (cell candidate) (39)
$c_{t}$ $=f_{t}\odot c_{t-1}+i_{t}\odot g_{t},\qquad h_{t}\;=\;o_{t}\odot\tanh(c_{t}).$ (40)
Define the diagonal gate matrices and slope matrices
$F_{t}=D(f_{t}),\quad I_{t}=D(i_{t}),\quad O_{t}=D(o_{t}),\quad G_{t}=D(g_{t}),$
$S^{i}_{t}=S^{\sigma}(a^{i}_{t}),\quad S^{f}_{t}=S^{\sigma}(a^{f}_{t}),\quad S^{o}_{t}=S^{\sigma}(a^{o}_{t}),\quad S^{g}_{t}=S^{\tanh}(a^{g}_{t}),$
and the cell-expression factors
$S_{t}=S^{\tanh}(c_{t})=D\!\bigl(1-\tanh^{2}(c_{t})\bigr),\qquad H_{t}=D\!\bigl(\tanh(c_{t})\bigr),\qquad E_{t}=D\!\bigl(o_{t}\odot(1-\tanh^{2}(c_{t}))\bigr)=O_{t}S_{t}.$
One-step Jacobian.
With $s_{t}=[h_{t};c_{t}]$, the Jacobian blocks are
$\frac{\partial c_{t}}{\partial c_{t-1}}$ $=F_{t},$ $\frac{\partial c_{t}}{\partial h_{t-1}}$ $=\underbrace{D(c_{t-1})\,S^{f}_{t}U_{f}}_{\text{via forget}}+\underbrace{I_{t}S^{g}_{t}U_{g}}_{\text{via candidate}}+\underbrace{G_{t}S^{i}_{t}U_{i}}_{\text{via input}}=:C_{t}^{(h)},$ (41)
$\frac{\partial h_{t}}{\partial c_{t-1}}$ $=E_{t}F_{t},$ $\frac{\partial h_{t}}{\partial h_{t-1}}$ $=H_{t}S^{o}_{t}U_{o}+E_{t}C_{t}^{(h)}.$ (42)
Collecting terms,
$J_{t}=\begin{bmatrix}H_{t}S^{o}_{t}U_{o}+E_{t}C_{t}^{(h)}&E_{t}F_{t}\\ C_{t}^{(h)}&F_{t}\end{bmatrix}\in\mathbb{R}^{2H\times 2H}.$ (43)
First-order expansion of the transport.
Decompose each one-step Jacobian in $\mathcal{M}_{t,\ell}$ as
$J_{t}=\mathcal{T}_{t}+\mathcal{R}_{t},$
where
$\mathcal{T}_{t}=\begin{bmatrix}0&E_{t}F_{t}\\ 0&F_{t}\end{bmatrix},\qquad\mathcal{R}_{t}=J_{t}-\mathcal{T}_{t}.$
Thus $\mathcal{T}_{t}$ contains the diagonal transport terms associated with cell-state retention ($F_{t}$) and final cell-to-hidden expression ($E_{t}$), while $\mathcal{R}_{t}$ collects the recurrently mixed corrections.
Applying the first-order expansion from Appendix A, we obtain:
$\mathcal{M}_{t,\ell}\;\approx\;\mathcal{T}_{t,\ell}\;+\;\sum_{p=\ell+1}^{t}\mathcal{T}_{t,p}\,\mathcal{R}_{p}\,\mathcal{T}_{p-1,\ell},\qquad\mathcal{T}_{a,b}=\prod_{j=b+1}^{a}\mathcal{T}_{j},\quad\mathcal{T}_{b,b}=I.$ (44)
Since the loss depends on $h_{t}$ (Section 3), only the top-right block of $\mathcal{M}_{t,\ell}$, mapping $c_{\ell}\mapsto h_{t}$, contributes directly to the gradient transport.
Zeroth-order contributions.
Because $\mathcal{T}_{t}$ is block upper triangular with a zero $h\!\to\!h$ block, the $(h,c)$ block of the transport $\mathcal{T}_{t:\ell}$ reduces to a single surviving path in which the signal remains in the cell state until time $t$:
$\mathcal{T}_{t,\ell}\,]_{h,c}=E_{t}\,\Phi_{t,\ell},\qquad\Phi_{a,b}=\prod_{j=b+1}^{a}F_{j},\quad\Phi_{b,b}=I.$
In particular, the $c\!\to\!h$ transition can occur only at the final time $t$. Extracting the diagonal gives the neuron-wise zeroth-order rates
$\gamma^{(0)}_{t,\ell}=\mathrm{diagvec}\!\bigl(E_{t}\Phi_{t,\ell}\bigr)\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=e_{t}^{(q)}\prod_{j=\ell+1}^{t}f_{j}^{(q)},\qquad e_{t}^{(q)}=o_{t}^{(q)}\left(1-\tanh^{2}(c_{t}^{(q)})\right).$ (45)
First-order diagonal correction.
From the first-order expansion (44), the contribution of a single recurrent-mixing insertion to the top-right block of the transport is
$\gamma^{(1)}_{t,\ell}=\sum_{p=\ell+1}^{t}\mathrm{diagvec}\!\Big([\,\mathcal{T}_{t,p}\,\mathcal{R}_{p}\,\mathcal{T}_{p-1,\ell}\,]_{h,c}\Big)\in\mathbb{R}^{H},$
where $\cdot\,]_{h,c}$ denotes the $H\times H$ block mapping $c_{\ell}\mapsto h_{t}$. Unlike the zeroth-order case, the presence of the recurrent correction $\mathcal{R}_{p}$ reintroduces $h\!\to\!h$ transport, allowing a single mixing event to occur at any intermediate time $p\in\{\ell+1,\dots,t\}$. This term therefore collects the diagonal components of the first-order Fréchet expansion, corresponding to neuron-wise self-couplings induced by a single insertion of the recurrent mixing blocks. Off-diagonal entries encode cross-neuron interactions and are intentionally discarded, since they do not admit a neuron-wise time-scale interpretation.
LSTM effective learning rates.
With global learning rate $\mu>0$, we define the per-lag, per-neuron effective learning rates
$\mu_{t,\ell}=\mu\bigl(\gamma^{(0)}_{t,\ell}+\gamma^{(1)}_{t,\ell}\bigr)\in\mathbb{R}^{H},\qquad\mu^{(q)}_{t,\ell}=\mu\bigl(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\bigr)$. (46)
These quantities act as scalar multipliers modulating the contribution of gradients originating at lag $(t-\ell)$ to parameter updates, providing a neuron-wise characterization of the effective learning dynamics induced by the LSTM gates. The transport factor appearing in Section 4 is
$\Gamma^{(q)}_{t,\ell}=\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}$.
B.3 GRU: one-step Jacobian and effective learning rates
Dynamics.
A GRU updates
$a^{z}_{t}$ $=W_{z}x_{t}+U_{z}h_{t-1}+b_{z},$ $z_{t}$ $=\sigma(a^{z}_{t}),$ (update/leak) (47)
$a^{r}_{t}$ $=W_{r}x_{t}+U_{r}h_{t-1}+b_{r},$ $r_{t}$ $=\sigma(a^{r}_{t}),$ (reset/filtering) (48)
$a^{g}_{t}$ $=W_{h}x_{t}+U_{h}(r_{t}\odot h_{t-1})+b_{h},$ $g_{t}$ $=\tanh(a^{g}_{t}),$ (candidate) (49)
$h_{t}$ $=(1-z_{t})\odot h_{t-1}+z_{t}\odot g_{t}.$ (50)
Let
$Z_{t}=D(z_{t}),\qquad R_{t}=D(r_{t}),\qquad G_{t}=D(g_{t}),$
$D_{h,t-1}=D(h_{t-1}),\qquad S^{z}_{t}=S^{\sigma}(a^{z}_{t}),\qquad S^{r}_{t}=S^{\sigma}(a^{r}_{t}),\qquad S^{g}_{t}=S^{\tanh}(a^{g}_{t}).$
One-step Jacobian.
Differentiating the leak and update paths gives
$\frac{\partial}{\partial h_{t-1}}\bigl[(1-z_{t})\odot h_{t-1}\bigr]$ $=(I-Z_{t})-D_{h,t-1}S^{z}_{t}U_{z},$ (51)
$\frac{\partial g_{t}}{\partial h_{t-1}}$ $=S^{g}_{t}U_{h}\bigl(R_{t}+D_{h,t-1}S^{r}_{t}U_{r}\bigr),$ (52)
$\frac{\partial}{\partial h_{t-1}}\bigl[z_{t}\odot g_{t}\bigr]$ $=G_{t}S^{z}_{t}U_{z}+Z_{t}\frac{\partial g_{t}}{\partial h_{t-1}}.$ (53)
Summing contributions yields
$J_{t}=(I-Z_{t})+(G_{t}-D_{h,t-1})S^{z}_{t}U_{z}+Z_{t}S^{g}_{t}U_{h}R_{t}+Z_{t}S^{g}_{t}U_{h}D_{h,t-1}S^{r}_{t}U_{r}.$ (54)
First-order expansion of the transport.
Decompose the one-step Jacobian as
$J_{t}=\mathcal{T}_{t}+\mathcal{R}_{t},$
where
$\mathcal{T}_{t}=I-Z_{t}$
is the diagonal retention operator induced by the update gate, and $\mathcal{R}_{t}$ collects all remaining recurrent mixing terms. For $\mathcal{M}_{t,\ell}$, the first-order expansion gives
$\mathcal{M}_{t,\ell}\;\approx\;\mathcal{T}_{t,\ell}\;+\;\sum_{p=\ell+1}^{t}\mathcal{T}_{t,p}\,\mathcal{R}_{p}\,\mathcal{T}_{p-1,\ell},\qquad\mathcal{T}_{a,b}=\prod_{j=b+1}^{a}\mathcal{T}_{j},\quad\mathcal{T}_{b,b}=I.$ (55)
Since $\mathcal{T}_{t}=I-Z_{t}$ is diagonal, we may equivalently write
$\mathcal{T}_{t,\ell}=\Phi_{t,\ell},\qquad\Phi_{a,b}=\prod_{j=b+1}^{a}(I-Z_{j}),\quad\Phi_{b,b}=I.$
Zeroth-order contributions.
The diagonal product $\Phi_{t,\ell}$ yields the primary update-gate retention envelope
$\gamma^{(0)}_{t,\ell}=\mathrm{diagvec}\!\bigl(\Phi_{t,\ell}\bigr)\in\mathbb{R}^{H},\qquad\gamma^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-z_{j}^{(q)}),$ (56)
which governs the zeroth-order neuron-wise time scales associated with the update gate.
In addition, the multiplicative action of the reset gate along the candidate pathway induces further diagonal attenuations that are useful to track explicitly. We therefore define the auxiliary diagonal pathwise envelopes
$\rho^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}r_{j}^{(q)},\qquad\eta^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-z_{j}^{(q)})\,r_{j}^{(q)},$ (57)
which capture pure reset attenuation and mixed update-reset attenuation, respectively. These quantities isolate the shrinkage associated with reset-controlled chains and are reported as complementary diagonal envelopes, rather than as additional terms in the strict $\mathcal{T}_{t}+\mathcal{R}_{t}$ product expansion. They are included in the effective learning rates for interpretability, as they capture additional multiplicative attenuation along the candidate pathway.
First-order diagonal correction.
The first-order diagonal correction associated with recurrent mixing is obtained by extracting the diagonal of the first-order term in (55):
$\gamma^{(1)}_{t,\ell}=\sum_{p=\ell+1}^{t}\mathrm{diagvec}\!\left(\mathcal{T}_{t,p}\,\mathcal{R}_{p}\,\mathcal{T}_{p-1,\ell}\right)=\sum_{p=\ell+1}^{t}\mathrm{diagvec}\!\left(\Phi_{t,p}\,\mathcal{R}_{p}\,\Phi_{p-1,\ell}\right)\in\mathbb{R}^{H}.$
This term collects the diagonal components arising from a single insertion of the recurrent mixing operator into an otherwise gate-diagonal transport. As in the LSTM case, off-diagonal components encode cross-neuron interactions and are intentionally discarded in order to preserve a neuron-wise time-scale interpretation.
GRU effective learning rates.
With global learning rate $\mu>0$, we define the per-lag, per-neuron effective learning rates as
$\mu^{(q)}_{t,\ell}=\mu\Bigl(\gamma^{(0,q)}_{t,\ell}+\rho^{(0,q)}_{t,\ell}+\eta^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\Bigr),\qquad q=1,\dots,H.$ (58)
These quantities provide a diagonal, neuron-wise summary of how update and reset gates jointly shape the effective learning dynamics across temporal lags. The transport factor appearing in Section 4 is
$\Gamma^{(q)}_{t,\ell}=\gamma^{(0,q)}_{t,\ell}+\rho^{(0,q)}_{t,\ell}+\eta^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}$,
so that the auxiliary envelopes $\rho^{(0)}_{t,\ell}$ and $\eta^{(0)}_{t,\ell}$ are included in $\Gamma$ and hence in the learnability envelope $f(\ell)$.
B.4 Baseline gated RNNs
We also consider the baseline gated RNN models introduced in [38]: a per-neuron (diagonal) gate, a shared (global) scalar gate, and a constant scalar gate. All follow the common update template
$h_{t}=(1-s_{t})\odot h_{t-1}+s_{t}\odot\tilde{h}_{t},\qquad\tilde{h}_{t}=\tanh(a^{h}_{t}),\qquad a^{h}_{t}=W_{h}x_{t}+U_{h}h_{t-1}+b_{h},$
but differ in how the gate $s_{t}$ is produced.
For these architectures, the neuron-wise transport factor takes the form
$\Gamma^{(q)}_{t,\ell}=\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}$,
where $\gamma^{(0,q)}_{t,\ell}$ denotes the zeroth-order gate-product envelope and $\gamma^{(1,q)}_{t,\ell}$ collects the diagonal first-order corrections arising from recurrent mixing terms.
Per-neuron gate (DiagGate)
The gate is computed coordinate-wise:
$a^{s}_{t}=W_{s}x_{t}+U_{s}h_{t-1}+b_{s},\qquad s_{t}=\sigma(a^{s}_{t})\in(0,1)^{H}.$ (59)
Define
$S_{t}=D(s_{t}),\qquad S^{s}_{t}=S^{\sigma}(a^{s}_{t}),\qquad S^{h}_{t}=S^{\tanh}(a^{h}_{t}).$
The exact one-step Jacobian $J_{t}=\partial h_{t}/\partial h_{t-1}\in\mathbb{R}^{H\times H}$ is
$J_{t}=\underbrace{(I-S_{t})}_{\text{leak}}+\underbrace{\bigl(D(\tilde{h}_{t})-D(h_{t-1})\bigr)S^{s}_{t}U_{s}}_{\text{gate sensitivity}}+\underbrace{S_{t}S^{h}_{t}U_{h}}_{\text{candidate path}}.$ (60)
The leak term $(I-S_{t})$ is diagonal; the remaining two terms contain $U_{s}$ and $U_{h}$ and are generally full rank, mixing information across neurons.
Effective learning rates.
Taking
$\mathcal{T}_{t}=I-S_{t},\qquad\mathcal{R}_{t}=J_{t}-\mathcal{T}_{t},$
the zeroth-order envelope is
$\gamma^{(0,q)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-s_{j,q}),$
yielding per-neuron effective learning rates
$\mu^{(q)}_{t,\ell}=\mu\bigl(\gamma^{(0,q)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\bigr)$,
where $\gamma^{(1,q)}_{t,\ell}$ arises from the diagonal of the first-order terms generated by the gate-sensitivity and candidate-path corrections in Eq. (60).
Shared global scalar gate (SharedGate)
The gate is a single scalar at each time step:
$a^{s}_{t}=w_{s}^{\top}x_{t}+u_{s}^{\top}h_{t-1}+b_{s},\qquad s_{t}=\sigma(a^{s}_{t})\in(0,1).$ (61)
The update is
$h_{t}=(1-s_{t})h_{t-1}+s_{t}\tilde{h}_{t},\qquad\tilde{h}_{t}=\tanh(a^{h}_{t}).$
Using
$\frac{\partial s_{t}}{\partial h_{t-1}}=s_{t}(1-s_{t})\,u_{s},$
the Jacobian is
$J_{t}=\underbrace{(1-s_{t})I}_{\text{leak}}+\underbrace{s_{t}S^{h}_{t}U_{h}}_{\text{candidate path}}+\underbrace{s_{t}(1-s_{t})(\tilde{h}_{t}-h_{t-1})u_{s}^{\top}}_{\text{rank-1 gate sensitivity}}.$ (62)
The last term is a rank-1 outer product and is the only source of cross-neuron coupling other than $U_{h}$.
Effective learning rates.
Taking
$\mathcal{T}_{t}=(1-s_{t})I,\qquad\mathcal{R}_{t}=J_{t}-\mathcal{T}_{t},$
the zeroth-order envelope is
$\gamma^{(0)}_{t,\ell}=\prod_{j=\ell+1}^{t}(1-s_{j}),$
which is identical for all neurons. Hence
$\mu^{(q)}_{t,\ell}=\mu\bigl(\gamma^{(0)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\bigr)$,
with neuron-dependent corrections $\gamma^{(1,q)}_{t,\ell}$ coming from the diagonal of the first-order terms generated by the candidate path and the rank-1 gate sensitivity in Eq. (62).
Constant scalar gate (ConstGate)
As a minimal baseline, the gate is fixed to a constant scalar $s\in(0,1)$, independent of $t$, inputs, and hidden state. The update is
$h_{t}=(1-s)h_{t-1}+s\tilde{h}_{t},\qquad\tilde{h}_{t}=\tanh(a^{h}_{t}),\qquad a^{h}_{t}=W_{h}x_{t}+U_{h}h_{t-1}+b_{h}.$ (63)
The Jacobian is
$J_{t}=\underbrace{(1-s)I}_{\text{fixed leak}}+\underbrace{sS^{h}_{t}U_{h}}_{\text{candidate path}}.$ (64)
Unlike DiagGate and SharedGate, there are no gate sensitivities, because $s$ is constant and not learned. The dynamics are therefore a rigid combination of identity and the candidate path, with a fixed leakage rate (1s)(1-s) applied to all neurons.
Effective learning rates.
Taking
$\mathcal{T}_{t}=(1-s)I,\qquad\mathcal{R}_{t}=J_{t}-\mathcal{T}_{t},$
the zeroth-order envelope is
$\gamma^{(0)}_{t,\ell}=(1-s)^{t-\ell},$
identical for all neurons. Thus
$\mu^{(q)}_{t,\ell}=\mu\bigl(\gamma^{(0)}_{t,\ell}+\gamma^{(1,q)}_{t,\ell}\bigr)$,
with $\gamma^{(1,q)}_{t,\ell}$ depending only on the diagonal of the first-order candidate-path correction.
B.5 Gates as implicit multi-rate optimizers
The effective learning rates derived above quantify how gates reweight BPTT contributions at lag $\ell$ on a per-neuron basis, even under a fixed global learning rate $\mu$ in SGD-like optimizers. Zeroth-order terms produce diagonal transport envelopes, while the first-order terms introduce anisotropy: recurrent matrices modulated by gate slopes mix coordinates and steer updates into privileged subspaces. As a result, gated RNNs act as implicit multi-rate optimizers: they induce heterogeneous, lag-dependent learning rates across neurons and selectively bias update directions. This interpretation of gates as multi-rate optimizers was originally developed in [38] for the baseline gated models of Section B.4 and extends naturally to the LSTM and GRU derivations above.
Appendix C Envelope validation
As discussed in Section 5.1, the envelope $f(\ell)$ defined in (8) relies on transport factors $\Gamma^{(q)}_{t,\ell}$ obtained from the first-order diagonal expansion of the Jacobian product $\mathcal{M}_{t,\ell}$. This expansion is what makes it possible to assign each hidden unit a lag-dependent rate $\mu^{(q)}_{t,\ell}$ with a transparent decomposition into gating and mixing contributions—a structure that cannot be read off the diagonal of $\mathcal{M}_{t,\ell}$ directly. This appendix verifies that the interpretability gained by the first-order expansion comes at no cost to the envelope’s decay profile: $f(\ell)$ preserves the decay trend of the exact-transport envelope $f_{\mathcal{M}}(\ell)=\sum_{q=1}^{H}|\Lambda^{(q)}[\mathcal{M}_{t,\ell}]_{qq}|$. We establish this with model-free linear statistics that assume no parametric form for the decay.
Setup.
[…]
Results.
Architecture Spearman $\rho$ Pearson $r$
ConstGate $1.000\pm 0.000$ $0.904\pm 0.046$
SharedGate $1.000\pm 0.000$ $0.940\pm 0.025$
DiagGate $1.000\pm 0.002$ $0.960\pm 0.045$
GRU $1.000\pm 0.000$ $0.986\pm 0.013$
LSTM $0.994\pm 0.007$ $0.986\pm 0.016$
All (75 runs) $\geq 0.979$ $\geq 0.82$
Table 1: Envelope validation across 5 architectures, 3 optimizers, and 5 seeds. Each row aggregates 15 runs (3 optimizers $\times$ 5 seeds); values are mean $\pm$ std. The bottom row reports the worst case across all 75 runs.
Spearman $\rho\geq 0.979$ across all 75 runs confirms that the first-order expansion never disrupts the lag ordering of the envelope. Pearson $r\geq 0.82$ (mean $0.96$) on the $\log_{10}$-transformed envelopes establishes that $\log f\approx a\,\log f_{\mathcal{M}}+b$, i.e. the two envelopes are related by a power-law rescaling $f\propto f_{\mathcal{M}}^{\,a}$. Such a rescaling preserves the monotonic and scaling structure on which the learnability analysis depends.
Appendix D Projected adaptive base rate via Rayleigh quotient
This appendix provides additional details on the construction of the neuron–lag adaptive base rate $\Lambda^{(q)}_{r,\ell}$ used in Section 4. The key idea is to extract the effective optimizer learning rate along the parameter-space direction associated with neuron $q$ at lag $\ell$.
Recall the parameter–state Jacobian
$B_{\ell}(\theta)=\frac{\partial h_{\ell}}{\partial\theta}\in\mathbb{R}^{H\times P},$
introduced in Eq. (4), where $H$ denotes the dimension of the hidden state and $P=\dim(\theta)$ the number of trainable parameters. Its $q$-th row
$B_{\ell}^{(q)}=\frac{\partial[h_{\ell}]_{q}}{\partial\theta}\in\mathbb{R}^{1\times P}$
identifies the direction in parameter space along which neuron $q$ is locally sensitive at time $\ell$. Denoting by
$b_{q}=(B_{\ell}^{(q)})^{\top}\in\mathbb{R}^{P},$
this vector specifies the principal parameter direction associated with neuron $q$ in the BPTT decomposition.
Under adaptive optimization (e.g. Adam or RMSprop), gradient updates are preconditioned by a diagonal matrix
$\Lambda_{r}=\mathrm{diag}(\lambda_{1,r},\dots,\lambda_{P,r}),$
so that the optimizer rescales each parameter coordinate independently.
Because $\Lambda_{r}$ acts anisotropically, the preconditioned sensitivity vector $\Lambda_{r}b_{q}$ is generally not collinear with $b_{q}$. Therefore, the optimizer does not simply rescale the neuron-specific direction but also distorts it in parameter space. To isolate the component of the update that remains aligned with the original sensitivity direction, we decompose
$\Lambda_{r}b_{q}=\Lambda^{(q)}_{r,\ell}\,b_{q}+r_{q},\qquad\langle b_{q},r_{q}\rangle=0,$
where the residual $r_{q}$ is orthogonal to $b_{q}$. The scalar coefficient $\Lambda^{(q)}_{r,\ell}$ therefore represents the magnitude of the orthogonal projection of the preconditioned vector $\Lambda_{r}b_{q}$ onto the one-dimensional subspace $\mathrm{span}(b_{q})$.
This projection coefficient is uniquely given by the Rayleigh quotient [21]
$\Lambda^{(q)}_{r,\ell}=\frac{\langle b_{q},\Lambda_{r}b_{q}\rangle}{\langle b_{q},b_{q}\rangle}=\frac{B_{\ell}^{(q)}\,\Lambda_{r}\,(B_{\ell}^{(q)})^{\top}}{B_{\ell}^{(q)}(B_{\ell}^{(q)})^{\top}}.$
Expanding the numerator yields
$\Lambda^{(q)}_{r,\ell}=\frac{\sum_{i=1}^{P}\lambda_{i,r}\,[B_{\ell}^{(q)}]_{i}^{2}}{\sum_{i=1}^{P}[B_{\ell}^{(q)}]_{i}^{2}}.$
Thus the effective base rate is a weighted average of the per-parameter adaptive learning rates, where the weights are determined by the squared parameter sensitivities of neuron $q$.
The orthogonal residual $r_{q}$ represents the component of the preconditioned update that changes the direction of the parameter step. While this residual affects the precise orientation of the update in parameter space, the scalar coefficient $\Lambda^{(q)}_{r,\ell}$ captures the amplification or attenuation of the update along the neuron’s principal sensitivity axis.
Finally, because $\Lambda^{(q)}_{r,\ell}$ is a convex combination of positive adaptive rates $\lambda_{i,r}$, it satisfies
$\min_{i}\lambda_{i,r}\leq\Lambda^{(q)}_{r,\ell}\leq\max_{i}\lambda_{i,r}.$ (65)
In particular $\Lambda^{(q)}_{r,\ell}>0$ whenever $B_{\ell}^{(q)}\neq 0$.
A useful consistency check is obtained in the case of plain SGD, where the optimizer preconditioner reduces to an isotropic scaling $\Lambda_{r}=\mu I$. Substituting this form into the Rayleigh quotient gives
$\Lambda^{(q)}_{r,\ell}=\frac{\langle b_{q},\mu Ib_{q}\rangle}{\langle b_{q},b_{q}\rangle}=\mu.$
Thus, the projected adaptive base rate collapses to the global learning rate $\mu$, and the generalized effective learning rates $\mu^{(q)}_{t,\ell}=\Lambda^{(q)}_{r,\ell}\Gamma^{(q)}_{t,\ell}$, where $\Gamma^{(q)}_{t,\ell}$ is the neuronwise state-space transport factor, recover the original SGD formulation $\mu^{(q)}_{t,\ell}=\mu\,\Gamma^{(q)}_{t,\ell}$.
Appendix E Noise floor of the empirical matched statistic
This appendix clarifies the assumptions underlying the noise model used in the manuscript and justifies the scaling of stochastic fluctuations in the empirical matched statistic. The key distinction is between exact finite-sample distributions and domain-of-attraction arguments, which suffice to characterize scaling behavior. Throughout, we rely on the linearity of the matched statistic and on standard properties of heavy-tailed random variables [42].
By construction, the empirical matched statistic at lag $\ell$ is an average over $N$ independent training sequences,
$\widehat{S}_{N}(\ell)=\frac{1}{N}\sum_{n=1}^{N}S^{(n)}_{t,\ell}$,
where $S^{(n)}_{t,\ell}$ denotes the statistic computed from the $n$th sequence. Independence across sequences implies that the summands are independent for fixed $(t,\ell)$. Empirical studies of stochastic gradients in deep neural networks consistently report heavy-tailed fluctuations, typically with tail index $\alpha\in(1,2)$. Accordingly, we assume that the centered per-sequence statistic $S^{(n)}_{t,\ell}-\mathbb{E}[S_{t,\ell}]$ belongs to the domain of attraction of a symmetric $\alpha$-stable law. This assumption does not require exact $\mathcal{S}\alpha\mathcal{S}$ behavior at finite sample sizes; it only requires regularly varying tails with index $\alpha$.
The matched statistic $S_{t,\ell}$ is a finite linear combination of gradient components (Eq. (13)). Linearity preserves the tail index of heavy-tailed variables, so $S_{t,\ell}-\mathbb{E}[S_{t,\ell}]$ also lies in the domain of attraction of a symmetric $\alpha$-stable law, with a lag-dependent scale parameter $\sigma_{\alpha}(\ell)$. For analytical convenience, we represent this domain of attraction by a canonical $\mathcal{S}\alpha\mathcal{S}(0,\sigma_{\alpha}(\ell))$ random variable. Stability under summation then implies
$\widehat{S}_{N}(\ell)-\mathbb{E}[S_{t,\ell}]\sim\mathcal{S}\alpha\mathcal{S}\!\left(0,\sigma_{\alpha}(\ell)N^{1/\alpha-1}\right)=\mathcal{S}\alpha\mathcal{S}\!\left(0,\frac{\sigma_{\alpha}(\ell)}{N^{\,1-1/\alpha}}\right).$
Note that the $N^{-1/\alpha}$ scaling applies to sums of $\alpha$-stable variables, whereas the empirical average introduces an additional factor $1/N$, yielding the rate $N^{-(1-1/\alpha)}$. Thus, the typical magnitude of stochastic fluctuations decays as $N^{-(1-1/\alpha)}$. For $\alpha=2$ this recovers the classical Gaussian rate $N^{-1/2}$, while for $\alpha<2$ concentration is slower due to heavy-tailed noise. This noise floor determines the detectability thresholds and sample-complexity bounds derived in the manuscript and underlies the compression of the learnability window in the heavy-tailed regime.
Appendix F LAN-based KL bound for averaged $\alpha$-stable location models
For completeness, we derive the KL lower bound used in Section 5.4.2 for the empirical matched statistic $\widehat{S}_{N}(\ell)$, which is defined as an average over $N$ independent training sequences. Unlike the classical LAN setting for product measures, the present analysis concerns a shrinking-scale $\alpha$-stable location family. The argument follows the general LAN framework for triangular arrays of statistical experiments with vanishing noise levels [52].
Statistical model and parameterization.
We consider the statistical experiment induced by the empirical matched statistic $\widehat{S}_{N}(\ell)$ at a fixed lag $\ell$, which is defined as an average over $N$ independent training sequences. As shown in Appendix E, under heavy-tailed gradient fluctuations with tail index $\alpha>1$, the centered statistic $\widehat{S}_{N}(\ell)-\mathbb{E}[S_{t,\ell}]$ exhibits stochastic fluctuations whose typical magnitude scales as $N^{-(1-1/\alpha)}$.
To formalize hypothesis testing at lag $\ell$, we model the distribution of the centered empirical statistic by a location family with shrinking scale. Specifically, let $Z$ be a symmetric $\alpha$-stable random variable with unit scale and characteristic exponent $\alpha>1$, and define the observation
$Y_{N}=\theta+s_{N}Z,\qquad s_{N}=\sigma_{\alpha}(\ell)\,N^{1/\alpha-1}.$ (66)
Here, $Y_{N}$ represents the random fluctuation of the empirical matched statistic, and $\theta\in\mathbb{R}$ is a location parameter encoding the presence or absence of a lag-$\ell$ dependency.
In the detection problem considered in the main text, $\theta$ corresponds to the mean shift induced by the effective learning rate alignment:
$\theta\in\left\{+\tfrac{1}{2}\Delta(\ell),\;-\tfrac{1}{2}\Delta(\ell)\right\},\qquad\Delta(\ell)=\overline{m}_{\mu}(\ell)\,f(\ell).$
Thus, the two hypotheses differ only by a small location perturbation whose magnitude is controlled by the effective learning-rate envelope.
For $\alpha>1$, symmetric $\alpha$-stable laws are absolutely continuous and admit a density. Denoting by $P_{\theta,N}$ the law of $Y_{N}$, its density is given by
$p_{\theta,N}(y)=\frac{1}{s_{N}}\,f_{0}\!\left(\frac{y-\theta}{s_{N}}\right),$ (67)
where $f_{0}$ is the density of a fixed symmetric $\alpha$-stable distribution with unit scale.
Triangular-array structure and relation to classical LAN.
Unlike the classical LAN framework, which considers $N$ i.i.d. observations from a fixed distribution, the present setting involves a triangular array of statistical experiments. This means that for each sample size $N$ we observe a single random variable $Y_{N}$, whose distribution depends on $N$ through the shrinking scale $s_{N}$.
Intuitively, this structure arises because $\widehat{S}_{N}(\ell)$ is itself an average over $N$ independent sequences: increasing $N$ does not provide more independent observations of a fixed noise model, but instead produces a more concentrated version of the same statistic. As $N$ grows, the noise scale $s_{N}$ tends to zero, yielding a sequence of location families with vanishing dispersion.
Such shrinking-scale location models are a standard object in asymptotic statistics [52] and admit local asymptotic normality under mild regularity conditions. In the present case, for $\alpha>1$, the family $\{P_{\theta,N}\}_{N\geq 1}$ satisfies a LAN expansion with normalization rate $r_{N}=s_{N}^{-1}$.
LAN for shrinking-scale location families.
For $\alpha>1$, symmetric $\alpha$-stable densities possess finite Fisher information for the location parameter [23, 34, 42, 52]. As $s_{N}\to 0$, the family $\{P_{\theta,N}\}_{N\geq 1}$ forms a triangular array of location models that is locally asymptotically normal with normalization rate
$r_{N}=\frac{1}{s_{N}}=\frac{N^{\,1-1/\alpha}}{\sigma_{\alpha}(\ell)}.$
Specifically, there exists a constant $I_{\alpha}>0$, depending only on the tail index $\alpha$, and a sequence of random variables $\Delta_{N}$ such that, for each fixed $h\in\mathbb{R}$,
$\log\frac{dP_{\theta+h/r_{N},N}}{dP_{\theta,N}}=h\,\Delta_{N}-\tfrac{1}{2}h^{2}I_{\alpha}+o_{P_{\theta,N}}(1),\qquad\Delta_{N}\xrightarrow{d}\mathcal{N}(0,I_{\alpha}),$ (68)
as $N\to\infty$. The constant $I_{\alpha}$ plays the role of an effective information coefficient in the local asymptotic expansion and does not require an explicit expression.
Taking expectations with respect to $P_{\theta,N}$ and using $\mathbb{E}[\Delta_{N}]=0$ yields the asymptotic expansion of the KL divergence [34]:
$D_{\mathrm{KL}}\!\big(P_{\theta+h/r_{N},N}\|P_{\theta,N}\big)=\tfrac{1}{2}h^{2}I_{\alpha}+o(1).$ (69)
Consequently, for all sufficiently large $N$, there exists a constant $c_{\alpha}>0$ such that
$D_{\mathrm{KL}}\!\big(P_{\theta+h/r_{N},N}\|P_{\theta,N}\big)\;\geq\;c_{\alpha}\,h^{2}.$ (70)
Application to the detection problem in the main text.
In the main text, the detection and non-detection hypotheses correspond to two symmetric location parameters separated by a fixed amount $\Delta(\ell)$:
$\theta_{1}=+\tfrac{1}{2}\Delta(\ell),\qquad\theta_{0}=-\tfrac{1}{2}\Delta(\ell),$
so that $\theta_{1}-\theta_{0}=\Delta(\ell)$.
The empirical matched statistic has a shrinking noise scale $s_{N}=\sigma_{\alpha}(\ell)N^{1/\alpha-1}$ (Appendix E). To express the fixed separation in the LAN framework, we measure it in units of this scale and define the normalized separation
$h_{N}=\frac{\theta_{1}-\theta_{0}}{s_{N}}=\frac{N^{\,1-1/\alpha}\,\Delta(\ell)}{\sigma_{\alpha}(\ell)}.$
Substituting $h=h_{N}$ into the LAN-based KL lower bound (70) yields
$D_{\mathrm{KL}}\!\left(P_{\text{det}}\|P_{\text{non}}\right)\;\geq\;c_{\alpha}\,h_{N}^{2}=c_{\alpha}\,\frac{N^{\,2(1-1/\alpha)}\,\Delta(\ell)^{2}}{\sigma_{\alpha}(\ell)^{2}},$ (71)
which is exactly the bound used in Eq. (17) of the main text.
Appendix G From KL divergence to mutual information and Fano bounds
This appendix provides a self-contained derivation of the mutual information lower bound and the Fano-type sample complexity inequality used in Sec. 5.4.2. The arguments rely only on standard relations between KL divergence, mutual information, and binary hypothesis testing [12].
Binary detection setup.
Fix a lag $\ell$. The hypotheses in Sec. 5.4.2 correspond to two $\alpha$–stable location models for the empirical matched statistic $S=\widehat{S}_{N}(\ell)$, obtained by averaging over $N$ independent training sequences. Introduce a binary label
$B\in\{0,1\},\qquad\mathbb{P}(B{=}1)=\mathbb{P}(B{=}0)=\tfrac{1}{2},$
where B=1B{=}1 selects the detection distribution $P_{\text{det}}$ and B=0B{=}0 selects the non-detection distribution $P_{\text{non}}$. The mutual information between $B$ and $S$ is defined as
$I(B;S)=D_{\mathrm{KL}}\!\big(P_{B,S}\,\|\,P_{B}P_{S}\big),$
where $P_{B,S}$ denotes the joint distribution and $P_{B}P_{S}$ the product of the marginals.
Mutual information via the mixture identity.
Because $B$ is equiprobable, the joint density factorizes as
$p_{B,S}(b,s)=p_{B}(b)\,p_{S\mid B}(s\mid b)=\tfrac{1}{2}\,p_{S\mid B}(s\mid b),$
with
$p_{S\mid B}(s\mid 1)=p_{\text{det}}(s),\qquad p_{S\mid B}(s\mid 0)=p_{\text{non}}(s).$
The marginal distribution of $S$ is the mixture
$M=\tfrac{1}{2}P_{\text{det}}+\tfrac{1}{2}P_{\text{non}}.$
Starting from the definition of mutual information,
$I(B;S)=\sum_{b=0}^{1}\int p_{B}(b)\,p_{S\mid B}(s\mid b)\log\frac{p_{S\mid B}(s\mid b)}{p_{S}(s)}\,dt,$
one obtains the mixture identity
$I(B;S)=\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{det}}\|M\big)+\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{non}}\|M\big).$ (72)
Since KL divergences are nonnegative,
$I(B;S)\;\geq\;\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{det}}\|M\big),$
and since mutual information cannot exceed the entropy of the binary label, $I(B;S)\leq\log 2$. Combining these observations yields the general lower bound
$I(B;S)\;\geq\;\min\!\left\{\log 2,\;\tfrac{1}{2}\,D_{\mathrm{KL}}\!\big(P_{\text{det}}\|M\big)\right\}.$ (73)
In the main text, the factor $\tfrac{1}{2}$ and the difference between $D_{\mathrm{KL}}(P_{\text{det}}\|M)$ and $D_{\mathrm{KL}}(P_{\text{det}}\|P_{\text{non}})$ are absorbed into the constant $c_{\alpha}$, since only the scaling with respect to $N$ and $\Delta(\ell)/\sigma_{\alpha}(\ell)$ is relevant.
Fano’s inequality and sample complexity.
Let $P_{e}$ denote the probability of incorrectly deciding whether a lag-$\ell$ signal is present based on the observation of $S$. For binary hypotheses with equal priors, Fano’s inequality [12] gives
$P_{e}\;\geq\;1-\frac{I(B;S)}{\log 2}.$
Thus, to ensure a target error probability $P_{e}\leq\epsilon<1/2$, it is sufficient that
$I(B;S)\;\geq\;\log 2\,(1-\epsilon).$
Substituting the mutual-information lower bound from Eq. (18) and solving for NN yields the sample complexity inequality
$N\;\geq\;\left(\frac{\sigma_{\alpha}(\ell)}{\sqrt{c_{\alpha}}\,\overline{m}_{\mu}(\ell)\,f(\ell)}\right)^{\!\frac{\alpha}{\alpha-1}}\!\!\Big(\log\tfrac{1}{2\epsilon}\Big)^{\frac{\alpha}{2(\alpha-1)}},$ (74)
which is the bound reported in Sec. 5.4.2.
Appendix H Proofs of Lemmas
Proof of Lemma 5.1
Proof.
By Section 4, the generalized effective learning rates factorize as
$\mu^{(q)}_{t,\ell}=\Lambda^{(q)}_{r,\ell}\Gamma^{(q)}_{t,\ell}$,
where $\Gamma^{(q)}_{t,\ell}$ is the neuronwise transport factor obtained from the diagonal terms of the first-order Jacobian-product expansion.
For all architectures considered in this paper, the zeroth-order transport terms are products of gate activations and bounded activation derivatives, hence they have the generic form
$\prod_{k=0}^{\ell-1}a^{(q)}_{t-k},\qquad a^{(q)}_{t-k}\in[0,1],$
and are therefore nonincreasing in $\ell$.
The same property holds for the first-order diagonal corrections. Indeed, each first-order term is obtained by inserting a single recurrent mixing contribution into an otherwise diagonal transport chain (Appendix A). Its diagonal entries are finite sums of terms consisting of an $\ell$-independent coefficient multiplied by products of gate factors in $[0,1]$. Increasing the lag can only append further multiplicative factors in $[0,1]$, and thus cannot increase the magnitude of these terms. Therefore
$|\Gamma^{(q)}_{t,\ell+1}|\leq|\Gamma^{(q)}_{t,\ell}|\qquad\text{for all }q.$
If the base rate is lag-independent, $\Lambda^{(q)}_{r,\ell}=\Lambda^{(q)}_{r}>0$, then
[…]
Proof of Lemma 5.2
Proof.
Starting from the per-lag sufficient sample size in Eq. (23),
$N(\ell)=\kappa_{\alpha,\epsilon}\,\left(\frac{\sigma_{\alpha}(\ell)}{\overline{m}_{\mu}(\ell)\,f(\ell)}\right)^{\!\kappa_{\alpha}},\qquad\kappa_{\alpha,\epsilon}=\frac{1}{c_{\alpha}^{\kappa_{\alpha}/2}}\Big(\log\tfrac{1}{2\epsilon}\Big)^{\kappa_{\alpha}/2},$
and applying the boundedness assumptions $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$ and $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$, we obtain
$\kappa_{\alpha,\epsilon}\left(\frac{c_{\sigma}}{C_{m}}\right)^{\!\kappa_{\alpha}}\,f(\ell)^{-\kappa_{\alpha}}\ \leq\ N(\ell)\ \leq\ \kappa_{\alpha,\epsilon}\left(\frac{C_{\sigma}}{c_{m}}\right)^{\!\kappa_{\alpha}}\,f(\ell)^{-\kappa_{\alpha}},$
which yields Eq. (24) with $c_{\star}=\kappa_{\alpha,\epsilon}(c_{\sigma}/C_{m})^{\kappa_{\alpha}}$ and $C_{\star}=\kappa_{\alpha,\epsilon}(C_{\sigma}/c_{m})^{\kappa_{\alpha}}$.
For the horizon bound, recall the detectability condition from Eq. (21):
$f(\ell)\geq\ \frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)}.$
As before, the constant factor $\frac{1}{\sqrt{c_{\alpha}}}\sqrt{\log\tfrac{1}{2\epsilon}}$ is independent of $\ell$ and $N$ and can be absorbed into the constants of the sandwich bound.
Bounding the right-hand side above and below using the assumed bounds on $\sigma_{\alpha}(\ell)$ and $\overline{m}_{\mu}(\ell)$ gives
$\frac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\ \leq\ \frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)}\ \leq\ \frac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}.$
The detectability condition is guaranteed whenever
$f(\ell)\geq\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}},$
and cannot hold unless
$f(\ell)\geq\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}.$
Since $f(\ell)$ is nonincreasing (Lemma 5.1), the admissible lags form an interval $0,f^{\leftarrow}(x)]$. Applying the generalized inverse $f^{\leftarrow}$ yields
$f^{\leftarrow}\!\Big(\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}\Big)\ \leq\ \mathcal{H}_{N}\ \leq\ f^{\leftarrow}\!\Big(\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\Big),$
which establishes the claimed sandwich bound. ∎
Generalized inverse.
The generalized inverse is used because $f(\ell)$ need not be strictly decreasing. By Lemma 5.1, $f(\ell)$ is nonincreasing, ensuring that $f^{\leftarrow}$ is well defined. The proof shows that the dependence on stochasticity and alignment enters only through the constants $(c_{\sigma},C_{\sigma},c_{m},C_{m})$ and the tail index $\alpha$ (via $\kappa_{\alpha}$); the functional dependence of $N(\ell)$ and $\mathcal{H}_{N}$ on the envelope $f(\ell)$ is dictated entirely by the power $f(\ell)^{-\kappa_{\alpha}}$.
On the boundedness assumptions.
The conditions $c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma}$ and $c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m}$ should be understood as mild regularity assumptions over the finite range of lags under consideration. Absent pathological degeneracies, such as vanishing alignment $\overline{m}_{\mu}(\ell)\to 0$ or diverging fluctuation scale $\sigma_{\alpha}(\ell)\to\infty$, these quantities remain finite and vary smoothly with $\ell$. Thus, the boundedness assumptions affect only constant factors in Eq. (24), while the asymptotic dependence of $N(\ell)$ and $\mathcal{H}_{N}$ on the envelope $f(\ell)$ is entirely captured by the exponent $\kappa_{\alpha}$.
Appendix I Asymptotic Scaling of the Learnability Window under $\alpha$-Stable Noise
This appendix provides the detailed derivations supporting the scaling relations summarized in Sec. 5.5 and established by Lemmas 5.15.2. Throughout, we use the corrected concentration exponent $\kappa_{\alpha}=\alpha/(\alpha-1)$, which governs the statistical behavior of empirical averages under $\alpha$-stable noise. Starting from the $\alpha$-stable model of Eq. (16) and the finite-sample bound of Eq. (19), we show how the slow $N^{1/\kappa_{\alpha}}$ concentration rate characteristic of heavy-tailed averages determines the asymptotic growth of the learnability window $\mathcal{H}_{N}$.
Per-lag sufficient sample size.
From Eq. (19), the minimal number of independent sequences required to detect a lag-$\ell$ dependency with error probability $P_{e}\!\leq\!\epsilon$ is
$N(\ell)\;=\;\kappa_{\alpha,\epsilon}\,\left(\frac{\sigma_{\alpha}(\ell)}{\overline{m}_{\mu}(\ell)f(\ell)}\right)^{\!\kappa_{\alpha}},\qquad\kappa_{\alpha,\epsilon}=\frac{1}{c_{\alpha}^{\kappa_{\alpha}/2}}\,\Big(\log\tfrac{1}{2\epsilon}\Big)^{\kappa_{\alpha}/2}.$ (75)
The constant $\kappa_{\alpha,\epsilon}$ depends only on the tail index $\alpha$ and the target detection error $\epsilon$, and can be absorbed into the asymptotic constants below. We assume the boundedness conditions of Lemma 5.2,
$c_{\sigma}\leq\sigma_{\alpha}(\ell)\leq C_{\sigma},\qquad c_{m}\leq\overline{m}_{\mu}(\ell)\leq C_{m},$
and recall that $\ell\mapsto f(\ell)$ is nonincreasing (Lemma 5.1).
Two-sided sandwich bound.
Substituting these bounds into (75) yields constants
$c_{\star}=\kappa_{\alpha,\epsilon}\Big(\tfrac{c_{\sigma}}{C_{m}}\Big)^{\!\kappa_{\alpha}},\qquad C_{\star}=\kappa_{\alpha,\epsilon}\Big(\tfrac{C_{\sigma}}{c_{m}}\Big)^{\!\kappa_{\alpha}},$
such that
$c_{\star}\,f(\ell)^{-\kappa_{\alpha}}\ \leq\ N(\ell)\ \leq\ C_{\star}\,f(\ell)^{-\kappa_{\alpha}}.$ (76)
Hence,
$N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$ (77)
gives the master proportionality relation between the minimum sample size and the envelope, formalizing that longer dependencies (smaller $f(\ell)$) require increasingly large sample sizes to detect under heavy-tailed noise.
Learnability window as a level set.
From the finite-sample detectability condition of Eq. (21),
$f(\ell)\ \geq\ \frac{\sigma_{\alpha}(\ell)}{N^{1/\kappa_{\alpha}}\,\overline{m}_{\mu}(\ell)},$
and using the same bounding arguments, we can express the learnability window $\mathcal{H}_{N}$ as a level set of $f(\ell)$:
$f^{\leftarrow}\!\Big(\tfrac{C_{\sigma}}{c_{m}}\,N^{-1/\kappa_{\alpha}}\Big)\ \leq\ \mathcal{H}_{N}\ \leq\ f^{\leftarrow}\!\Big(\tfrac{c_{\sigma}}{C_{m}}\,N^{-1/\kappa_{\alpha}}\Big),$ (78)
where $f^{\leftarrow}(y)=\sup\{\ell\geq 1:f(\ell)\geq y\}$ denotes the generalized inverse. By Lemma 5.1, $f(\ell)$ is nonincreasing, ensuring that the generalized inverse $f^{\leftarrow}$ is well-defined. This reproduces the sandwich bound in Eq. (25) of the main text.
Asymptotic regimes.
Considering the master proportionality in Eq. 77, we now invert (76) (equivalently (78)) for three canonical decay laws of the envelope $f(\ell)$.
(i) Logarithmic decay.
If $f(\ell)\asymp c/\log(1+\ell)$ with $c>0$, then
$N(\ell)\asymp[\log(1+\ell)]^{\kappa_{\alpha}}\quad\Longrightarrow\quad\log(1+\ell)\asymp N^{1/\kappa_{\alpha}}\quad\Longrightarrow\quad\mathcal{H}_{N}\asymp\exp(\kappa N^{1/\kappa_{\alpha}})-1.$
(ii) Polynomial (algebraic) decay.
If $f(\ell)\asymp c\,\ell^{-\beta}$ with $\beta>0$, then
$N(\ell)\asymp\ell^{\kappa_{\alpha}\beta}\quad\Longrightarrow\quad\mathcal{H}_{N}\asymp N^{1/(\kappa_{\alpha}\beta)}.$
(iii) Exponential (geometric) decay.
If $f(\ell)\asymp c\,\lambda^{\ell}$ with $\lambda\in(0,1)$, then
$N(\ell)\asymp\lambda^{-\kappa_{\alpha}\ell}\quad\Longrightarrow\quad\mathcal{H}_{N}\asymp\frac{\log N}{\kappa_{\alpha}\,\log(1/\lambda)}.$
All asymptotic forms hold up to multiplicative constants inherited from $c_{\sigma}$, $C_{\sigma}$, $c_{m}$, $C_{m}$, and $\kappa_{\alpha,\epsilon}$. As $\alpha$ decreases, heavier tails increase $\kappa_{\alpha}$ and slow the $N^{1/\kappa_{\alpha}}$ concentration of empirical averages, uniformly compressing the learnability window $\mathcal{H}_{N}$ across all decay regimes.
Appendix J Optimizer dependence of envelope scaling and time-scale spectra
[…]
Our framework predicts that architectural design and optimization dynamics jointly shape the coupled state–parameter system and may bias training toward different regimes. However, it does not require that a single regime be universally realized across optimizers, tasks, or training configurations. In particular, the emergence of broad mixtures of effective time scales depends on the interaction between architecture, optimizer, sequence length, task structure, and training hyperparameters. Different configurations may therefore lead to qualitatively different envelope behavior even under the same optimizer. The results below document one consistent realization of the diagnostic pipeline under SGD-like dynamics, without asserting optimizer-specific regimes.
Plain SGD.
[…]
Consistently, the time-scale spectra remain bounded (Fig. 7), supporting a finite-memory exponential regime under plain SGD in the present setting.
[…]
Figure 7: Plain SGD: time-scale spectra $\tau_{q}$. All architectures exhibit bounded support, consistent with exponential envelope decay.
Noise statistics: plain SGD.
[…]
The noise scale $\hat{\sigma}_{\alpha}(\ell)$ decays rapidly with lag (Fig. 8(b)), more strongly than observed under AdamW. This accelerated attenuation of gradient fluctuations is consistent with the finite-memory exponential envelope regime identified above.
Overall, plain SGD does not eliminate heavy-tailed behavior entirely, but it exhibits weaker and more rapidly attenuating tail effects compared to the AdamW setting. The concentration properties therefore remain compatible with the theoretical framework, albeit with reduced heavy-tailed persistence.
[…]
(b) Noise scale $\hat{\sigma}_{\alpha}(\ell)$ under plain SGD (log scale).
Figure 8: Noise diagnostics under plain SGD.
SGD with momentum.
[…]
The corresponding time-scale spectra remain bounded (Fig. 10). Although characteristic time scales shift relative to plain SGD, no heavy-tailed or unbounded $\tau_{q}$ distribution emerges. Thus, SGD with momentum also resides in a finite-memory exponential regime in the present setting.
[…]
Figure 9: SGD with momentum: envelope scaling diagnostics.
Figure 10: SGD with momentum: unit-level time-scale spectra $\tau_{q}$.
Noise statistics: SGD with momentum.
The estimated tail indices remain heavy-tailed and well separated from the Gaussian limit (Fig. 11(a)). The noise scale $\hat{\sigma}_{\alpha}(\ell)$ decays smoothly with lag (Fig. 11(b)) without optimizer-induced regime transitions.
Therefore, although momentum alters effective decay rates and shifts characteristic time scales, it does not modify the underlying heavy-tailed noise class in the present configuration. Observed differences in envelope behavior primarily reflect changes in effective time-scale geometry rather than transitions in statistical concentration properties.
[…]
(b) Noise scale $\hat{\sigma}_{\alpha}(\ell)$ under momentum (log scale).
Figure 11: Noise diagnostics under SGD with momentum.
Appendix K Additional time-scale spectrum diagnostics
Tail-emphasis representation.
To further examine the slow end of the time-scale spectra, Fig. 12 reports the complementary cumulative distribution functions (CCDFs) on log–log axes. This representation emphasizes differences in intermediate and large time scales. ConstGate and SharedGate exhibit sharply truncated spectra, with negligible mass beyond a narrow range of $\tau$. In contrast, DiagGate, GRU, and LSTM display extended intermediate ranges over which the CCDF decays slowly, followed by a clear attenuation at the largest observed $\tau$. While these intermediate regimes may appear approximately linear over limited ranges, the eventual curvature confirms that the distributions are not asymptotically scale-invariant, but instead reflect a mixed multi-scale regime with an effective finite-size cut-off.
Figure 12: Log–log representation of the CCDF of neuronwise time scales $\tau_{q}$. DiagGate, GRU, and LSTM exhibit extended intermediate tails with eventual attenuation, whereas ConstGate and SharedGate are tightly concentrated.
Per-model distributions.
For completeness, Fig. 13 shows the empirical probability density functions (histograms) of $\tau_{q}$ for each model. ConstGate and SharedGate produce narrowly concentrated spectra, consistent with their homogeneous gating geometry. DiagGate exhibits a visibly heterogeneous mixture, with a majority of short time scales and a smaller subpopulation of substantially longer ones. GRU and LSTM display broader and more gradually decaying spectra, indicating richer mixtures of effective memory depths. These density-level diagnostics reinforce the CCDF-based conclusions reported in Sec. 6 and confirm that envelope geometry and learnability horizon are tightly linked to the dispersion of neuronwise time scales.
(a) ConstGate
(b) SharedGate
(c) DiagGate
(d) GRU
(e) LSTM
Figure 13: Empirical probability density functions of neuronwise time scales $\tau_{q}$. DiagGate, GRU, and LSTM exhibit visibly broader and more heterogeneous spectra than ConstGate and SharedGate.
Appendix L Code availability
The code used to reproduce the experiments reported in this paper is available at
https://github.com/lorenzolivi/learnability
The repository includes scripts and instructions for reproducing the results.